### v1

# Simulations of gravitational collapse in null coordinates IV: evolving through the event horizon, with an application to the spherical charged scalar field

Abstract
[…]
I Introduction
I.1 The problem posed by evolving through an event horizon in null coordinates
In this series of papers, starting with axinull_formulation , we have been developing numerical simulations of gravitational collapse on null cones emanating from a regular centre: in the present paper, we examine formulations that allow us to evolve from regular data through the black hole event horizon. After this Introduction, we will restrict to spherical symmetry in 3+1 spacetime dimensions, but the ideas generalise straightforwardly, and will be applied to non-spherical critical collapse elsewhere.
In Paper II of the series axinull_critscalar , we investigated the threshold of black hole formation for a massless scalar field coupled to general relativity in axisymmetry, adapting the coordinates to the expected self-similar critical solution. While, at moderate non-sphericity, this allowed us to fine-tune to the threshold of black hole formation to machine precision, without any need for adaptive mesh-refinement, we could not reach non-sphericities as high as others had reached before us. In hindsight, this was due both to our choice of radial gauge (how the radial coordinate $x$ labels points on each null cone generator) and our choice of formulation (which of the Einstein equations are used in the code and which are not). For our notation for the metric, see already Eq. (1) below.
Our gauge choice in Paper II was to make the area radius $R$ a function $R(u,x)$ of retarded time $u$ and radial coordinate $x$ only, but not the two angular coordinates: we called this local shifted Bondi (lsB) gauge, as it is a generalisation of Bondi gauge $R=x$. We noted that near the event horizon, in highly non-spherical spacetimes, this was not a good gauge choice, as it caused the radial shift $B$ required to keep $R$ independent of the angles to diverge.
Separately, we chose the formulation where the metric coefficient $G$ is obtained by solving the Raychaudhuri equation along each outgoing null coordinate line. This formulation requires the area radius $R$ to be strictly increasing along each null cone generator, that is $R_{,x}>0$. In spherical symmetry, this still allows our time slices (outgoing null cones) to approach the event horizon asymptotically. Beyond spherical symmetry, there will be retarded times $u_{0}$ and $u_{1}$ such that the full event horizon is seen only after $u=u_{1}$, but in a formulation that requires $R_{,x}>0$ the code must stop already at $u=u_{0}$.
The combination of lsB gauge and the formulation where $G$ is constrained meant in practice that the simulations of highly non-spherical spacetimes near the black-hole treshold had to stop before they could be classified as either collapsing or dispersing.
This paper presents a second formulation, where the Raychaudhuri equation is solved for $R$, rather than for $G$. In the literature, this approach has been used only together with the affine radial gauge $G=1$. Here we evolve $G$, which allows us to use an arbitrary radial shift $B$, and in particular one adapted to critical collapse. We also present a third formulation where both $R$ and $G$ are evolved, and the Raychaudhuri equation is not used. In the literature, this has been used only with double null gauge $B=0$, but $B$ can in fact be chosen freely. Here we point out that all three formulations can be combined with any radial gauge choice. We present extended tests of these three formulations in spherical symmetry, with charged scalar field matter.
The gauge problem in critical collapse mentioned above does not arise in spherical symmetry, and will be addressed when we return to axisymmetry elsewhere.
I.2 Formulations and radial gauges in null coordinates
We now give a more precise statement of what we just said about formulations, but without restricting the spacetime dimension or symmetry. (In the remainder of the paper, we will restrict to spherical symmetry in 3+1 dimensions.) We remind the reader that we consider line elements of the form
$ds^{2}$ $=$ $-2G\,du\,(dx+B\,du)$ (1)
$+R^{2}\gamma_{ij}(d\theta^{i}+\beta^{i}\,du)(d\theta^{j}+\beta^{j}\,du).$
(The quantity $B$ replaces the metric coefficient $H:=-g_{uu}$ that we used in Papers I–III. This is just for convenience of notation.) The spacetime dimension is $n+2$, with the $\theta^{i}$ coordinates on an $n$-dimensional manifold, typically $S^{n}$. From this form of $g_{\mu\nu}$ it follows that $g^{uu}=g^{ui}=0$, and $g^{ux}=-G^{-1}$. Hence the surfaces of constant $u$ are null surfaces, and their affinely parameterised generators have tangent vector
$U^{a}:=-\nabla^{a}u=G^{-1}(\partial_{x})^{a}.$ (2)
We must have $G>0$ for the metric to be regular. $R$ is defined by restricting $\gamma_{ij}$ to have a fixed determinant (such as its value in Minkowski spacetime). The remaining gauge freedom in this ansatz for the line element is to re-label $u$ as $u\to u^{\prime}(u)$, and to relabel the radial coordinate $x$ as $x\to x^{\prime}(u,x,\theta^{i})$. We call a choice of coordinate $x$ a radial gauge.
The expansion of the congruence of null generators can be expressed in terms of $R$ as
$\rho_{+}={d\ln R\over d\lambda},$ (3)
where we have defined
${d\over d\lambda}:=G^{-1}\partial_{x}.$ (4)
(See also Paper I axinull_formulation for the full definition of $\rho_{+}$.) The quantity
$\lambda(u,x,\theta^{i}):=\int_{0}^{x}G(u,x^{\prime},\theta^{i})\,dx^{\prime}$ (5)
is a preferred affine parameter along the null generators. (We shall refer to it as “the” affine parameter).
The $xx$-component of the trace-reversed Einstein equations can be written as
$R_{,xx}-{G_{,x}\over G}R_{,x}+KR=0,$ (6)
or as
$\left(\ln{G\over R_{,x}}\right)_{,x}={KR\over R_{,x}},$ (7)
or as
${d^{2}R\over d\lambda^{2}}+{K\over G^{2}}R=0,$ (8)
or as the Raychaudhuri equation
${d\rho_{+}\over d\lambda}=-\rho_{+}^{2}-{K\over G^{2}},$ (9)
where the term $K$ contains shear and matter terms, and is non-negative in vacuum or with matter obeying the dominant energy condition.
From (8), with $K/G\geq 0$, we see that once $dR/d\lambda$ has become negative, $R$ reaches zero in finite affine parameter. Equivalently, from (9) we see that once $\rho_{+}$ has become negative, it blows up in finite affine parameter. We can now make the above statement precise: In a non-spherical collapse spacetime, generically there will be $u_{0}<u_{1}$ such that for any coordinate null cone with $u_{0}<u<u_{1}$, the expansion $\rho_{+}$ remains positive on some of its generators, while on others it becomes negative and then blows up. This could be because it has reached a curvature singularity, or because the null cone has developed a caustic.
We now come to the role of the Raychaudhuri equation in numerical relativity on null cones. In its form (6) and with $G$ given, it can be solved as a second-order ordinary differential equation (ODE) in $x$ for $R$. In its form (7) and with $R$ given, it can be solved as a first-order ODE for $G$. We also have wave equation-like evolution equations $G_{,ux}=$... and $R_{,ux}=$... for both $G$ and $R$. This gives rise to three formulations, each of which admits an arbitrary radial gauge:
  • evolve $R$ and constrain $G$ (evolve-$R$ formulation, from now on eR);
  • evolve $G$ and constrain $R$ (evolve-$G$ formulation, from now on eG);
  • evolve $R$ and $G$ freely and impose the constraint between them only at $u=0$ (free-evolution formulation, from now on fe).
The eR formulation requires $R_{,x}>0$, because we divide by $R_{,x}$ in (7), and so it breaks down once any part of a coordinate null cone intersects the black hole region. The other two formulations allow us to evolve further. It is this observation that motivates us to consider eG and fe in the present paper.
After the radial gauge has been fixed in the initial data at $u=0$, typically by specifying either $R$ or $G$ as a function of $(x,\theta^{i})$, it is controlled incrementally by the choice of the quantity $B$. The prototype incremental gauge choice is double-null gauge $B=0$, which makes $x$ an ingoing null coordinate.
In addition to the three formulations above, there are two particular combinations of formulation and gauge that remove the need for one evolution equation, and which we feel should be characterised as formulations:
  • set $R=x$ and constrain $G$ (Bondi formulation);
  • set $G=1$ and constrain $R$ (affine formulation).
In both these formulations, all time evolution equations (equations containing $u$-derivatives) correspond to physical (matter and gravitational wave) degrees of freedom, so they are “maximally constrained” in the language of numerical relativity. By contrast, eR and eG have one gauge evolution equation, and fe has two. Table 1 summarises the five formulations we have just discussed. The observation that the choices of formulation and gauge are independent is one of the main points of the present paper.
Formulation Initial gauge Evolved gauge Restricted to Gauge evolution equations
Bondi $R=x$ $R=x$ $R_{,x}>0$ 0
affine $G=1$ $G=1$ 0
eR any $R$ or $G$ any $B$ $R_{,x}>0$ 1
eG any $R$ or $G$ any $B$ 1
free evolution any $R$ or $G$ any $B$ 2
Table 1: Summary of the five formulations discussed in Sec. I.2. “Initial gauge” denotes the freedom to label points on $u=0$ by the coordinate $x$, so “any RR” is short for “any choice of $R(0,x,\theta^{i})$” and similarly for “any GG”. “Evolved gauge” denotes how the coordinate $x$ is propagated to $u>0$, and “any BB” is short for “any choice of $B(u,x,\theta^{i})$”. The restriction to $R_{,x}>0$ arises when the Raychaudhuri equation is solved for $G$ given $R$, and means that we cannot evolve through an event horizon. We have suppressed the function arguments in the table so that it applies for any spatial symmetry or none, and any spacetime dimension. “Gauge evolution equations” denotes the number of equations containing $u$-derivatives in the formulation, minus the number of physical degrees of freedom (where in the absence of symmetry we count vacuum gravity as two degrees of freedom, a scalar field as one, etc).
We are aware of the following combinations of formulation and gauge choice used for numerical time evolution in the literature. We hope that the list of such combinations is complete, but the list of references is only intended to be indicative:
In our own previous work, we have explored the following additional combinations:
I.3 Plan of this paper
In the following, we will restrict to spherical symmetry in 3+1 dimensions, with charged scalar field matter, but the ideas illustrated here are easily generalised to less symmetry, other spacetime dimensions, and to vacuum or other matter.
We present an implementation of all three formulations set out above in one code, and test them in two settings. In the first of these, we evolve on outgoing null cones that emanate from a regular central worldline. To fix the radial gauge in the initial data, we either set $R=x$ or $G=1$ at $u=0$. We continue in “shifted double-null gauge”, an incremental gauge that is adapted to critical collapse in spherical symmetry (see Sec. III.C of Paper I for a definition and PortoGundlach2022 for a previous application). As a physics application, we compute the fine-structure of the mass and charge scaling laws in critical collapse, which has not been computed before. In the second setting the left boundary $x=0$ is not a regular centre but an ingoing null cone, so that the computational domain is a null rectangle. We continue the initial gauge choice of $R_{x,}=1$ or $G=1$ as double-null gauge $B=0$.
In both settings, we present evidence of second-order convergence with resolution, both within each formulation and between the three pairs of formulations. With a regular centre, the convergence is not quite pointwise at the first few grid points, but it is pointwise elsewhere, and is second-order in any norm. By contrast, on the null rectangle, where we do not need an expansion, the convergence is cleanly pointwise.
We devise a simple singularity excision scheme, and present evidence that it works in the eG and free evolution formulations (eR cannot evolve through the apparent horizon).
In Sec. II we give the field equations for the spherical Einstein-charged-scalar system, in Sec. III we present our solution algorithm, and in Sec. IV our numerical methods. Sec. V presents tests on a domain with a regular centre, including convergence tests, critical collapse, and singularity excision, and Sec. VI presents tests on a null rectangle domain, including convergence tests, collapse, and singularity excision. We conclude in Sec. VII.
II Field equations
II.1 Field equations in covariant form
The matter is a complex scalar $\phi=:\psi+i\chi$ field coupled to electromagnetism. The action, in units where $c=G_{N}=1$, is
$S=\int\left({R\over{16\pi}}-{{1\over 2}}D_{a}\phi(D^{a}\phi)^{*}-{1\over{16\pi}}F_{ab}F^{ab}\right)\sqrt{-g}\,d^{4}x,$ (10)
where a star denotes the complex conjugate, the field strength tensor $F$ in terms of the potential $A$ is
$F_{ab}:=\nabla_{a}A_{b}-\nabla_{b}A_{a},$ (11)
and we have introduced the charge-covariant derivative
$D_{a}:=\nabla_{a}+iqA_{a}.$ (12)
Our electromagnetic convention is similar to Gauss units, and our scalar field convention reduces to the standard one for a real scalar field $\psi$ if $\chi=0$. These are also the conventions of HodPiran97 . The convention of GellesPretorius25 differs by the absence of the factor $1/2$ in the scalar field term. The convention of MurReaTan13 differs by factors of $16\pi$ in front of $R$ and $4\pi$ in front of $F_{ab}F^{ab}$.
$\phi$ obeys the wave equation
$D^{a}D_{a}\phi=0.$ (13)
The Maxwell equations are
$\nabla_{b}F^{ab}=4\pi j^{a}$}, (14)
where the charge current is
$j_{a}$ $=$ $-{iq\over 2}(D_{a}\phi)^{*}\phi+c.c.$ (15)
$=$ $q(\chi\nabla_{a}\psi-\psi\nabla_{a}\chi)-q^{2}A_{a}(\psi^{2}+\chi^{2}).$ (16)
It is conserved, $\nabla_{a}j^{a}=0$, because $F^{ab}$ is antisymmetric. The trace-reversed Einstein equations are
$E_{ab}:=R_{ab}-8\pi S_{ab}=0,$ (17)
where the trace-reversed stress-energy tensor is
$S_{ab}$ $=$ ${{1\over 2}}\left(D_{a}\phi(D_{b}\phi)^{*}+(D_{a}\phi)^{*}D_{b}\phi\right)$ (18)
$+{{1\over 4\pi}}\left(F_{ac}{F_{b}}^{c}-{1\over 4}g_{ab}F_{cd}F^{cd}\right).$
The equations admit the gauge freedom $\phi\to e^{iq\alpha}\phi$, $A_{a}\to\nabla_{a}\alpha$ for an arbitrary scalar function $\alpha$, leaving $F_{ab}$ and $D_{a}\phi$, and hence $j^{a}$ and $T_{ab}$, invariant.
II.2 Field equations in spherical symmetry in null coordinates
We consider the general metric in spherical symmetry with one null coordinate $u$,
$ds^{2}=-2G\,du(dx+B\,du)+R^{2}\,d\Omega^{2}.$ (19)
where $G$, $B$ and $R$ are functions of $(u,x)$, and $d\Omega^{2}:=d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}$. The surfaces of constant $u$ are outgoing null cones. The choice of coordinate $x$ along the null cone generators (the radial gauge) is not yet fixed.
Starting from an arbitrary electromagnetic gauge, the particular gauge
$A_{x}(u,x)=0,\qquad A_{u}(u,0)=0$ (20)
can be achieved by the gauge transformation with
$\alpha(u,x)=\int_{0}^{u}A_{u}(u^{\prime},0)\,du^{\prime}-\int_{0}^{x}A_{x}(u,x^{\prime})\,dx^{\prime}.$ (21)
From now on, we work exclusively in this gauge, and for conciseness we rename $A_{u}$ to $A$. Note that $A_{x}=0$ can be written geometrically as $U^{a}A_{a}=0$.
We define the ingoing null vector field
$\Xi:=\partial_{u}-B\partial_{x},$ (22)
which is normalised relatively to $U^{a}$ as $U^{a}\Xi_{a}=-1$. For the scalar field, we also introduce
$\hat{\Xi}\phi:=(\Xi+iqA)\phi,$ (23)
with real and imaginary parts
$\hat{\Xi}\psi$ $:=$ ${\rm Re}\,\hat{\Xi}\phi=\Xi\psi-qA\chi,$ (24)
$\hat{\Xi}\chi$ $:=$ ${\rm Im}\,\hat{\Xi}\phi=\Xi\chi+qA\psi.$ (25)
The two components of the Maxwell equations can be written in terms of the local charge function
$Q:={R^{2}A_{,x}\over G}.$ (26)
The Maxwell equations become
$Q_{,x}$ $=$ ${4\pi}qR^{2}(\psi\chi_{,x}-\chi\psi_{,x}),$ (27)
$\Xi Q$ $=$ ${4\pi}qR^{2}(\chi\hat{\Xi}\psi-\psi\hat{\Xi}\chi).$ (28)
Hence we have a conserved charge current $\nabla_{a}j^{a}=0$ with
$j^{u}={1\over{4\pi}GR^{2}}Q_{,x},\quad j^{x}=-{1\over{4\pi}GR^{2}}Q_{,u},$ (29)
where an overall constant has been fixed from (14), and we have used $\sqrt{-g}=GR^{2}\sin\theta$.
We can write the Einstein and Maxwell equations in terms of $\Xi$, $\hat{\Xi}$ and $Q$, without any explicit appearance of $A$, $B$ and $\partial_{u}$, with the exception of the Einstein equation $E_{ux}$. However, we still need to find $B$ in order to recover $\partial_{u}$ from $\Xi$, and we need to find $A$ to recover $\Xi$ from $\hat{\Xi}$. We find $A$ by integrating (26), written as
$A_{,x}={GQ\over R^{2}}.$ (30)
In spherical symmetry, the Hawking mass $M$ of a symmetry sphere is a scalar on the reduced spacetime, and is then also called the Misner-Sharp mass, or the Kodama mass. It is given by
$M:={CR\over 2},$ (31)
where we define the Hawking compactness $C$ of a symmetry sphere (equivalent to a point in the reduced spacetime) as
$C:=1-|\nabla R|^{2}.$ (32)
In our coordinates, $C$ is given by
$C(u,x)=1+{2R_{,x}\Xi R\over G}.$ (33)
As $R\geq 0$ and $G>0$, $\rho_{+}$ has the same sign as $R_{,x}$. Hence a symmetry sphere is trapped if and only if $R_{,x}<0$, and marginally trapped if and only if $R_{,x}=0$. (By trapped, we always mean outer-trapped.)
The derivatives of the Misner-Sharp mass are
$M_{,x}$ $=$ $-4\pi{R^{2}\Xi R\over G}(\psi_{,x}^{2}+\chi_{,x}^{2})+{Q^{2}R_{,x}\over 2R^{2}},$ (34)
$\Xi M$ $=$ $-4\pi{R^{2}R_{,x}\over G}(\hat{\Xi}\psi^{2}+\hat{\Xi}\chi^{2})+{Q^{2}\Xi R\over 2R^{2}}.$ (35)
Hence there is a conserved mass curent analogous to (29).
We also define the augmented mass
${\cal M}:=M+{Q^{2}\over 2R}.$ (36)
[…]
${\cal M}_{,x}$ $=$ $-4\pi{R^{2}\Xi R\over G}(\psi_{,x}^{2}+\chi_{,x}^{2})$ (37)
$+{4\pi}qQR(\psi\chi_{,x}-\chi\psi_{,x}),$
$\Xi{\cal M}$ $=$ $-4\pi{R^{2}R_{,x}\over G}(\hat{\Xi}\psi^{2}+\hat{\Xi}\chi^{2})$ (38)
$+{4\pi}qQR(\chi\hat{\Xi}\psi-\psi\hat{\Xi}\chi).$
Hence there is another conserved mass curent analogous to (29).
[…]
In spherical symmetry, only the components $E_{uu}$, $E_{ux}$, $E_{xx}$ and $E_{+}:=E_{\theta\theta}=E_{\varphi\varphi}/\sin^{2}\theta$ of the trace-reversed Einstein equations are algebraically independent. We can write $E_{xx}=0$ as
$\left(\ln{G\over R_{,x}}\right)_{,x}=4\pi{R(\psi_{,x}^{2}+\chi_{,x}^{2})\over R_{,x}},$ (39)
which can be integrated for $G$, given $R$, $\psi$ and $\chi$, but only if $R_{,x}>0$. Alternatively, we can write it as
$R_{,xx}-{G_{,x}\over G}R_{,x}+4\pi R(\psi_{,x}^{2}+\chi_{,x}^{2})=0,$ (40)
which can be solved as a second-order linear ODE for $R$, given $G$, $\psi$ and $\chi$. There is then no restriction on the sign of $R_{,x}$. More geometrically, $E_{xx}=0$ can be written as
$UUR+4\pi[(U\psi)^{2}+(U\chi)^{2}]R=0.$ (41)
It is the Raychaudhuri equation for the generators of the coordinate null cones.
We can write $E_{+}=0$ and the real and imaginary parts of the complex wave equation as
$(R\Xi R)_{,x}$ $=$ ${-{G\over 2}\left(1-{Q^{2}\over R^{2}}\right)},$ (42)
$(R\hat{\Xi}\psi)_{,x}$ $=$ $-(\Xi R)\psi_{,x}-{qQG\chi\over 2R},$ (43)
$(R\hat{\Xi}\chi)_{,x}$ $=$ $-(\Xi R)\chi_{,x}+{qQG\psi\over 2R}.$ (44)
$E_{ux}=0$ can be written as
${\cal H}_{,x}$ $=$ $-{G+2R_{,x}\Xi R\over R^{2}}+{{2}GQ^{2}\over R^{4}}$ (45)
$+8\pi(\chi_{,x}\hat{\Xi}\chi+\psi_{,x}\hat{\Xi}\psi)$
$=$ ${-{2G{\cal M}\over R^{3}}+{3GQ^{2}\over R^{4}}+8\pi(...)},$ (46)
where
${\cal H}:=B_{,x}-\Xi\ln G$ (47)
is the only place where $B$ appears in the Einstein equations, other than in the combination $\Xi$. $E_{ux}$ is a combination of derivatives of the other Einstein equations, and is therefore sometimes called the “redundant” equation, but in the eG and fe formulations we use it to obtain $\Xi\ln G$, given an arbitrary $B$, while in the affine formulation we use it to obtain $B$, given $G=1$.
[…]
The remaining Einstein equation $E_{uu}=0$ can be written as
$\Xi\Xi R+{\cal H}\Xi R+4\pi R(\hat{\Xi}\psi^{2}+\hat{\Xi}\chi^{2})=0.$ (48)
It is the Raychaudhuri equation on the ingoing null geodesics. We do not use it in our time evolution scheme with a regular centre, but it acts as a on constraint on the data we can impose on a left boundary that is ingoing null. In analogy to (41), we can write it more geometrically as
$X\!XR+4\pi[(\hat{X}\psi)^{2}+(\hat{X}\chi)^{2}]R=0,$ (49)
where
$X:=\bar{G}^{-1}\Xi,\qquad\hat{X}:=\bar{G}^{-1}\hat{\Xi}$. (50)
$X$ is tangent to the affinely parameterised ingoing null geodesics (as $U$ is to the outgoing ones), and $\bar{G}$ is defined by
$\Xi\ln\bar{G}=-{\cal H},$ (51)
or equivalently by
$\Xi\ln\left({\bar{G}\over G}\right)=-B_{,x}.$ (52)
In double-null gauge $B=0$, we have $\bar{G}=G$ and $X=G^{-1}\partial_{u}$, in analogy to $U=G^{-1}\partial_{x}$ (the latter holds in any radial gauge).
In initial data on an outgoing null cone with regular centre, we always have $\Xi R<0$ and hence $XR<0$ on our initial data. From (49) we see that $XXR\leq 0$, so $XR<0$ and hence $\Xi R<0$ on the entire domain of dependendence of these initial data. Hence, from (33), a symmetry sphere is trapped if and only if $C>1$ and marginally trapped if and only if $C=1$. This is not obvious from (42), whose right-hand side becomes negative when $R<|Q|$, and so, as we solve (42) but not (49) in algorithm, $\Xi R<0$ may be violated by numerical error.
The choice of coordinate $x$ for $u>0$ is controlled incrementally by the choice of $B$. This plays a role similar to the radial shift vector component in the ADM formulation. To see this, note that if we make a transformation from double-null coordinates $(u,v)$ to general null coordinates $(u,x)$, parameterised by $v=v(u,x)$, then
$-2\bar{G}\,du\,dv$ $=$ $-2G\,du(dx+B\,du),$ (53)
$G$ $=$ $\bar{G}v_{,x},$ (54)
$v_{,u}$ $=$ $Bv_{,x},$ (55)
regardless of spacetime dimension or symmetry.
[…]
$T:=\max_{\text{whole spacetime}}|(D_{a}\phi)^{*}D^{a}\phi|$ (56)
as an additional diagnostic. Note that $\max|{R_{a}}^{a}|=8\pi T$. This gives us a curvature quantity that scales on the dispersion side of the collapse threshold.
III Algorithm
III.1 Initial data and gauge choice
To set initial data, we specify $\psi(x)$ and $\chi(x)$ freely. We then either initialise $G=1$ and solve (40) to initialise $R$, or we initialise $R=x$ and solve (39) to initialise $G$. These are physically inequivalent data for the same functions $\psi(x)$ and $\chi(x)$, as they correspond to different geometrically defined data functions $\psi(R)$ and $\chi(R)$, or equivalently $\psi(\lambda)$ and $\chi(\lambda)$.
In Bondi gauge $R=x$, the identity $R_{,u}=0$ determines $B$ as
$B_{\text{Bondi}}=-{\Xi R\over R_{,x}}.$ (57)
All our gauge choices keep $R=0$ at $x=0$, and so $R_{,u}=0$ at the centre. With $u$ defined to be proper time there, we have
$B(u,0)=B_{\text{Bondi}}(u,0)={1\over 2R_{,x}(u,0)},$ (58)
see also Eq. (115) below.
In affine gauge $G=1$, the identity $\Xi G=0$ determines $B$ as
$B_{\text{affine}}=B_{\text{Bondi}}(u,0)+\int_{0}^{x}{\cal H}\,dx^{\prime},$ (59)
where we have used (58).
For critical collapse in spherical symmetry, we use a radial gauge where $B$ is a linear function of $x$, fixed to coincide with Bondi gauge at $x=0$ (so that $R=0$ remains at $x=0$) and to vanish at $x=x_{0}$ (so that $x=x_{0}$ is an ingoing null cone). This gives
$B_{\text{sdn}}(u,x)=B_{\text{Bondi}}(u,0)\left(1-{x\over x_{0}}\right).$ (60)
It is easy to see from (55) that if $v=x$ on $u=0$ and $B$ is proportional to $x$, then $v$ remains linear in $x$, with $v=x_{0}$ at $x=x_{0}$ and $R=0$ at $x=0$, for all $u\geq 0$. This simple linear relation between $v$ and $x$ motivates the name “shifted double-null” (sdn) gauge.
III.2 Boundary conditions at $x=0$
Where $R=x=0$ is a regular central worldline, $x=0$ is a coordinate singularity (and we omit it from the numerical grid), so we cannot impose simple boundary conditions there. To start up the integration of the hiercharchy equations, we expand in powers of $x$ instead.
At a given $u$, we fit the evolved quantities $\psi$, $\chi$, $R/x$ and/or $G$ either to linear functions of $x$, such as $\psi\simeq\psi_{(0)}+\psi_{(1)}x$, or to quadratic ones, such as $\psi\simeq\psi_{(0)}+\psi_{(1)}x+\psi_{(2)}x^{2}$. (The notation here is compatible with our notation $\psi_{l(n)}(u)$ in Papers I and II, where $l$ denotes the spherical harmonic and $n$ the power of $x$.)
$R_{(1)}$, $G_{(0)}$ and ${\cal H}_{(0)}$ are not determined by the equations. However, we show in Appendix B that at a regular centre, and with $u$ proper time at the centre, they obey $G_{(0)}=R_{(1)}$, $B_{(0)}=1/(2R_{(1)})$ and ${\cal H}_{(0)}=0$.
The full expressions are given in Appendix C.
On the other hand, if $x=0$ is a timelike or null inner boundary, we can specify boundary conditions for the hierarchy equations there, subject to constraints. For the case where $x=0$ is an ingoing null cone, we discuss this in detail in Appendix E.
III.3 Time evolution algorithm
We use the “method of lines” framework, where one discretises separately in space and time. We break up each sub-timestep into the following steps, where Steps 1 and 3 depends on the formulation.
Step 0: Assume that on a null cone of constant $u$, the functions $\psi(x)$, $\chi(x)$, and $\ln G(x)$ and/or $R(x)$ are given. When $G$ and $R$ are both given, they must obey the constraint (40). With a regular centre, we fit these functions to power-series expansions in $x$ about $x=0$.
Step 1a: In the eR or Bondi formulation, we solve the Raychaudhuri equation in its form (39) for $G$ by integration, and set the fitting coefficients of $G$ consistently.
Step 1b: Alternatively, in the eG or affine formulation, we solve the Raychaudhuri equation in its form (40) as a homogeneous linear second-order ODE for $R$, and we set the fitting coefficients of $R$ consistently.
Step 1c: Alternatively, in free evolution, we do nothing. From numerical error, both in the evolution of $R$ and $G$, and in the fitting, the fitting coefficients of $R$ and $G$ will not be exactly compatible with the Raychaudhuri equation, so we need to make an arbitrary choice which set of fitting coefficients we choose for expanding the solution of the hierarchy equations at the centre. The choice made in Appendix C works.
[…]
Step 3a: In the affine formulation, we compute $B$ from (59).
Step 3b: Alternatively, in the Bondi formulation, $B$ is given explicitly by (57).
Step 3c: Alternatively, in the other formulations we set $B$ freely. For critical collapse in spherical symmetry, we use sdn gauge (60). On a null rectangle we use double-null gauge $B=0$.
Step 4: We find the $u$-derivatives of $\psi$, $\chi$, $R$ and/or $\ln G$ from their $\Xi$-derivatives and $x$-derivatives, using (22). The $x$-derivatives are upwinded depending on the sign of $B$. At $x=0$ this makes the upwind derivative a right derivative, and at $x_{\text{max}}>x_{0}$ a left derivative.
Step 5: Using the method of lines, we evolve $\psi$, $\chi$, and one or both of $R$ and $\ln G$, from $u$ to $u+\Delta u$. This returns us to Step 0.
IV Numerical methods
IV.1 Discretization
With $x=R=0$ a regular centre, we use an equally spaced centered grid $x_{i}=i\Delta x$, $i=1...N_{x}$, so that $x=0$ is not on the grid. With $x=0$ an ingoing null cone, we use $x_{i}=i\Delta x$, $i=0,1...N_{x}$.
The ODEs in $x$ can be solved by integration using an integrating factor. They can therefore be written as $g_{,x}=F(f,f_{,x})$, where the $f$ are already known and $g$ is to be determined. We use the explicit integration scheme
$g_{i}\simeq g_{i-1}+F\left(f_{i-{1\over 2}},(f_{,x})_{i-{1\over 2}}\right)\,\Delta x,$ (61)
with
$f_{i-{1\over 2}}$ $\simeq$ ${f_{i}+f_{i-1}\over 2},$ (62)
$(f_{,x})_{i-{1\over 2}}$ $\simeq$ ${f_{i}-f_{i-1}\over\Delta x}.$ (63)
This scheme is by construction second-order accurate, and is causal and explicit in the sense that the vector of variables $g_{i}$ depends only on $g_{j}$ with $j<i$ and $f_{j}$ with $j\leq i$: information propagates along a null cone of constant $u$ only towards increasing $x$.
To solve the Raychaudhuri equation (40) for $R$, we write it in the first-order form
$V_{,x}$ $=$ $-{K\over G}\,R,$ (64)
$R_{,x}$ $=$ $G\,V,$ (65)
with
$V$ $:=$ ${R_{,x}\over G},$ (66)
$K$ $:=$ $4\pi(\psi_{,x}^{2}+\chi_{,x}^{2}).$ (67)
When we solve (64,65) for $(V,R)$, the coefficients $G$ and $K$ are already known.
$K$ is naturally given at midpoints, but $G$ needs to be interpolated there using (62). We discretize this system, using a 2-point stencil, to second-order accuracy as
${R_{i}-R_{i-1}\over\Delta x}$ $=$ $G_{i-{1\over 2}}{V_{i}+V_{i-1}\over 2},$ (68)
${V_{i}-V_{i-1}\over\Delta x}$ $=$ $-{K_{i-{1\over 2}}\over G_{i-{1\over 2}}}{R_{i}+R_{i-1}\over 2}.$ (69)
This can be solved for $R_{i}$ and $V_{i}$, giving us the explicit scheme
$\left(\begin{array}[]{c}R_{i}\\ V_{i}\\ \end{array}\right)={1\over 1+{{\cal K}}}\left(\begin{array}[]{cc}1-{{\cal K}}&{\cal G}\\ -{4{{\cal K}}\over{\cal G}}&1-{{\cal K}}\\ \end{array}\right)\left(\begin{array}[]{c}R_{i-1}\\ V_{i-1}\\ \end{array}\right),$ (70)
where
${\cal K}$ $:=$ $K_{i-{1\over 2}}{\Delta x^{2}\over 4},$ (71)
${\cal G}$ $:=$ $G_{i-{1\over 2}}\Delta x.$ (72)
The two complex eigenvalues of this linear map have absolute value 1 for any ${{\cal K}}\geq 0$, so that the method is unconditionally stable. The method is exact for $K=0$. It has no amplitude error when $K_{,x}=G_{,x}=0$, in the sense that then $(KR^{2}+G^{2}V^{2})_{,x}=0$ in the discretisation, as it is in the continuum.
$V(x)$ changes sign at most once (from positive to negative) over the $x$-domain. This means that $K\lesssim x_{\text{max}}^{-2}$, and so ${{\cal K}}\lesssim N_{x}^{-2}\ll 1$. Hence we can expect the numerical error to be reasonably small.
For comparison, we also use the second or fourth-order Adams-Bashforth method. These need the vector of right-hand sides $F_{i}$ at grid points. We discrete derivatives on grid points as
$(f_{,x})_{i}\simeq{f_{i+1}-f_{i-1}\over 2\Delta x}$ (73)
where this is defined, and the standard one-sided three-point stencils at the left and right boundaries $i=1$ and $i=N_{x}$. This is also second-order accurate but not causal.
When we solve (39) for $G$, we first find $\ln V$, where $V:=R_{,x}/G$, on grid points by integration. (In Papers I-III, we used the variables $g:=1/V$ and $\gamma:=\ln g$. We have replaced them with $V$ as they are defined only for $R_{,x}>0$.) We then find $V$ at grid points and hence at mid-points, and hence $G=R_{,x}/V$ at mid-points in a causal and second-order accurate way, given that $R_{,x}$ is known to second order and causally at mid-points.
We linearly interpolate $G$ from the mid-points to grid points (and we linearly extrapolate to the grid points $i=1$ and $i=N_{x}$). This is second-order accurate but with an additional error from the interpolation, and it is no longer causal. However, we need $G$ at grid points only for output and to initialise $\ln G$ from initial data in Bondi gauge. The non-causality does not matter for output and initialisation. To evaluate the compactness $C$ when $G$ is constrained, we use $V$ on grid points.
As $G$ varies rapidly in $u$ and $x$, but is constrained to be strictly positive, we evolve $\ln G$ rather than $G$. We find $G$ at grid points by exponentiating, and hence on mid-points by linear interpolation. To evaluate the compactness $C$ when $G$ is evolved, we us $R_{,x}$ on gridpoints, which is not causal. Again we do not see an adverse effect.
In the evolution equations
$f_{,u}=\Xi f+Bf_{,x}$ (74)
where $f$ stands for $\psi$, $\chi$, $R$ and $\ln G$, we discretize the transport term $f_{,x}$ by second-order upwinding, depending on the sign of $B$. If our solution method for the hierarchy equations is also causal, as defined above, our complete numerical method is strictly causal, in the sense that all grid points are computed only from grid points in their causal past. In particular, with $B<0$ at $i=N_{x}$ and $i=N_{x}-1$ , the outer boundary is treated numerically like any other point.
In any gauge where $R=x$ stays at $x=0$ we need to evaluate $R_{,x}$ at the centre to impose (58). Given that the upwinded transport term $BR_{,x}$ should cancel $\Xi R$ to give us $R_{,u}=0$ at the centre, we compute $R_{,x}$ at the centre as the second-order upwind derivative
$R_{(1)}:=R_{,x}(u,0)\simeq{2R_{1}-{1\over 2}R_{2}\over\Delta x}.$ (75)
Other choices result in an instability at the centre.
We solve the $f_{,u}=F[f]$ evolution equations using a Runge-Kutta ODE solver (typically second-order) in $u$, and we solve the ODEs in $x$ that give $F[f]$ at each sub-timestep.
IV.2 Quantities that scale as a power of $x$ at the centre
The quantity $Q$ is obtained by integration of (27). From its definition, the quantity $Q/R^{3}$, representing an average charge density, is expected to be smooth, and to be non-zero at the centre. Numerically, this is of course not guaranteed. We can improve regularity by discretizing
$\int_{x_{i-1}}^{x_{i}}f\,R^{2}\,dx\simeq{1-\epsilon+{\epsilon^{2}\over 3}\over 1-\epsilon+{\epsilon^{2}\over 4}}\ f_{i-1/2}\ (R_{i-1/2})^{2}\ \Delta x,$ (76)
where $\epsilon:=1/i$. This is exact when $R/x$ and $f$ are constant. Applying this to the computation of $Q$ does not make a big difference for $Q$ itself, but it does for $A$, which is the integral of $Q/R^{2}$. We apply the same correction factor in the integral for $\tilde{M}$, to correct for a naive midpoint value of $Q/R\sim x^{2}$.
For a similar reason, whereever we need to evaluate $M/R^{3}$, $Q/R^{2}$ and similar quantities, in the mid-point rule we take the average of $M/R^{3}$ at $x_{i}$ and $x_{i-1}$, rather than averaging $M$ and $R$.
[…]
IV.3 Choice of time step
In the spherical and axisymmetric scalar critical collapse simulations of Paper II we used the timestep $\Delta u=\Delta u_{1}$, where
$\Delta u_{1}:=C_{1}\min_{x}\left|{R_{,x}\over\Xi R}\right|\,\Delta x.$ (77)
(In axisymmetry, the minimum was taken also over the angular coordinate $y$). As is clear from the formula, this restricts the change of $R$ with $\Delta u$ in the ingoing null direction, compared to the change with $\Delta x$ in the outgoing null direction. $\Delta u=\Delta u_{1}$ with $C_{1}=0.1$ was used in all simulations in Paper II (although this was not stated there). These simulations used the eR formulation, and were stopped as soon as $\max_{x}C\geq 0.99$. As $C=1$ implies $R_{,x}=0$, $\Delta u_{1}\to 0$ as $\max_{x}C\to 1_{-}$.
The hierarchy equations are in effect ODEs in $x$, while the time update takes the form (74). Hence a necessary condition for the stability of our explicit scheme is that these advection equations obey the Courant-Friedrichs-Levy (CFL) condition. We guarantee this by restricting the timestep to
$\Delta u_{2}:=C_{2}\left(\max_{x}\left|B\right|\right)^{-1}\,\Delta x$ (78)
for some $C_{2}\lesssim 1$. In all our gauges $R_{,u}=0$ at $R=0$, and so (77) and (78) are equivalent at the centre.
Since Paper II, we have realised that using $\Delta u=\Delta u_{1}$ may be an unnecessarily small timestep. We have found that $\Delta u=\Delta u_{2}$ with $C_{1}=0.1$ also leads to stable evolutions up to close to the apparent horizon, but with fewer timesteps overall.
As the apparent horizon is approached, $\Delta u_{1}\to 0$, and so this timestep does not allow us to evolve through the horizon. The alternative timestep $\Delta u_{2}$ remains finite, and the evolution remains stable, but as the solution changes very rapidly with $u$ at the horizon, $\Delta u_{2}$ gives large errors there. (In Appendix D we remind the reader that $R$ on constant-$u$ time slices deviates from the horizon radius exponentially in $x$, and this in turn means that $\Xi R$ near the horizon grows exponentially in $x$.)
This problem was already noted by Burko and Ori in BurOri97 in simulations in spherical symmetry in double null coordinates. They were able to address it in an elegant way: their code advances an ingoing null slice in $v$, so that, for computational purposes, $v$ is the “time” coordinate, not $u$. The necessary mesh refinement in $u$ near the horizon is then made adaptively, by splitting grid cells in $u$ near $u=u_{h}$, increasingly finely as $v$ increases.
In our framework, we could in principle introduce adaptive mesh refinement to better resolve the event horizon, but this goes against our main motivation for using null coordinates in the first place, which is to resolve type-II critical collapse down to arbitrarily small scales by adapting the coordinates to self-similarity, without the need for explicit adaptive mesh refinement.
Hence to maintain accuracy through $u=h_{h}$, we just reduce the global timestep near the horizon. This works because near the horizon the solution changes rapidly in $u$ but not in $x$. In type-II critical collapse, most of the computation is in the approximately self-similar phase, and so is not affected by this relatively inefficient method of mesh refinement. After some experimentation we have come up with the timestep
$\Delta u_{4}=\min(\Delta u_{2},\Delta u_{3}),$ (79)
where
$\Delta u_{3}:=C_{3}{\max_{x}R\over\max_{x}|R_{,u}|},$ (80)
with $C_{3}=0.001$. Note $\Delta u_{3}$ and hence $\Delta u_{4}$ is not proportional to $\Delta x$, in contrast to $\Delta u_{1}$ and $\Delta u_{2}$.
If we want to evolve through a horizon, where $R_{,x}=0$, we cannot use the numerical time step $\Delta u_{1}$. In double-null gauge $B=0$, the time step $\Delta u_{2}$ is not defined. The time steps $\Delta u_{3}$ and $\Delta u_{4}$ cannot be used for convergence testing as $\Delta u_{3}$ is independent of $\Delta x$. As an alternative to all these, for example if we want to test convergence of a solution where $R_{,x}=0$ occurs, we can also use the simple timestep
$\Delta u_{0}:=C_{0}\Delta x.$ (81)
IV.4 Singularity excision
As no influence propagates from smaller to larger $x$ in our code for all $x$ with $B<0$, and we are in spherical symmetry, we can use a crude type of similarity excision where we excise gridpoints $i\geq i_{\text{exc}}$, where $i_{\text{exc}}$ is the first grid point where either $R_{i}<0$ or $R_{i}<R_{i-1}/2$. (The second criterion makes sure that the first non-excised point is not too close to $R=0$.) It is essential that we re-determine $i_{\text{exc}}$ after each time we have solved (40) for $R$, that is, in all sub-timesteps. Obviously $i_{\text{exc}}$ is not allowed to decrease from one sub-timestep to the next, or one timestep to the next.
Once excision starts, the excision radius $x_{\text{exc}}(u):=i_{\text{exc}}(u)\Delta x$ decreases quickly to zero. Once it is smaller than $x_{0}$, our upwind differencing is still to the right at the outer boundary. To avoid this, we adjust $x_{0}$ so that it remains the same fraction of the excision radius $x_{\text{exc}}$ as it was originally to $x_{\text{max}}$.
V Tests with a regular centre
V.1 Charged initial data
We set initial data of the form
$\psi(0,x)+i\chi(0,x)=p{\cal A}\,e^{-\left({x-x_{c}\over d}\right)^{2}+i\omega(x-x_{c})},$ (82)
so that for $\omega=0$ the scalar field is purely real, and hence for small $\omega$ we expect the charge to be proportional to $\omega$. With a charged scalar field, $q$ sets an inverse length scale in the field equations, so $qx_{c}$, $qd$ and $q/\omega$ are all relevant dimensionless parameters. This also means that changing $q\to\lambda q$ is equivalent to $x_{c}\to x_{c}/\lambda$, $d\to d/\lambda$, and similarly for all other length scales in the initial data and $\omega\to\lambda\omega$.
V.2 Convergence tests
For convergence testing, we have focused on initial data (82) with $x_{c}=0.5$, $d=0.5$, $\omega=2.0$, and the coupling constant $q=1/\sqrt{4\pi}$. We complete the initial data either with $G=1$, and then set $p{\cal A}=0.15$, or with $R=x$, and then set $p{\cal A}=0.2$. These amplitudes are chosen so that $\max_{x}C\simeq 0.5$ in the initial data at $x\simeq 1.2$.
Our numerical domain is determined by the choices $x_{\text{max}}=4.0$ and $x_{0}=3$. Both initial data sets collapse promptly. In the evolution $\max_{x}C\simeq 0.9$ is reached at $u\simeq 0.6$ at $x\simeq 1.2$, and the event horizon is crossed soon after. In these two solutions, the time step does not need to be reduced very much to resolve the horizon (in contrast to near-critical collapse). For each set of initial data, we evolve with the eR, eG and fe formulations. We thus have six combinations of initial data and formulation to consider.
We evolve with $N_{x}=100\cdot 2^{n}$ grid points for n=08n=0...8. We choose the timestep $\Delta u_{4}$ with parameters $C_{2}=0.1$, $C_{3}=10^{-3}$. We output all variables at intervals $\Delta u=0.1$, adjusting the last timestep before each output time down in order to hit it exactly. We evolve up to horizon formation, and compare resolutions up to the last output time before that. Our convergence testing methods are summarised in Appendix G. Near the horizon, the timestep $\Delta u_{4}$ can be smaller than the timestep $\Delta u_{2}$ that is by definition proportional to $\Delta x$. This does not seem to affect second-order convergence, presumably because the finite-differencing error in $x$ dominates over that in $u$.
We have used convergence testing to experiment with fitting expansions near $x=0$ of the evolved functions $\psi$, $\chi$, $R/x$ and/or $G$ to different numbers of points, and to linear or quadratic functions, and independently with expanding the solutions of the hierarchy equations near $x=0$ to first or second order, and for different numbers of points before our midpoint integration rule is applied. For this experimentation, we have found it helpful to first focus on the initial data, solving only the hierarchy equations, and to make decisions for one hierarchy equation at a time, in the order in which they are solved.
We eventually decided to expand some variables to second order, in the sense that we expand as far as we can using quadratic fits of $\psi$, $\chi$, $R/x$ and/or $G$. We then use the results of those quadratic fits, even where we expand only to first order. With that settled, we experimented with the order of expansion and the number of points to expand to, for each variable, in the order in which they are computed. We found that for all variables, expanding only to $i=1$, and finding $i=2,3$... by integration, was optimal. For most variables, expanding only to first order was optimal. These are the variables for which the errors $h^{2}f_{2}$ and $h^{3}f_{3}$ defined in (141) below are finite at $x=0$.
When expanding $R\hat{\Xi}\psi$ and $R\hat{\Xi}\chi$ to first order, we found that the errors $f_{2}$ and $f_{3}$ of $\hat{\Xi}\psi$ and $\hat{\Xi}\chi$ diverge as $x^{-1}$. This is not in itself a problem, as $h^{2}f_{2}(x_{i})\simeq h^{2}/(ih)=h/i$, so even at $i=1$ the error goes down as $h$. However, we find that by expanding $R\hat{\Xi}\psi$ and $R\hat{\Xi}\psi$ to second order, we can essentially remove the $h^{2}f_{2}$ error, so that the error is dominated by $h^{3}f_{3}$. This reduces the actual error at $i=1$ by two or three orders of magnitude already at $h=0.04$, and by more at higher resolution.
We have also output the quantities $Q/R^{3}$, $M/R^{3}$ and $\tilde{M}/R^{3}$ and tested their convergence. Again, they converge to second order, but the errors go as $x^{-1}$, $x^{-2}$ and $x^{-1}$, respectively, at the centre. This is why we use $\tilde{M}/R^{3}$ where we need to evaluate $M/R^{3}$. The correction factors in integration $Q$ and $\tilde{M}$ in Eq. (76) are essential for reducing the error in $Q/R^{3}$ and $\tilde{M}/R^{3}$ from $x^{-2}$ to $x^{-1}$.
[…]
With the methods discussed above, we achieve, roughly speaking, second-order pointwise convergence almost until the horizon is crossed. (We lose convergence at the horizon simply because some variables change very rapidly across the horizon, while the $u$-value of the horizon depends on the resolution.)
To state our convergence results more precisely, it is useful to split all variables into two groups. The variables in the first group, including $R$, $Q$, AA, $M$, $\tilde{M}$, and $V$ (in eR and eG) show clear pointwise second-order convergence including at the centre.
[…]
We note in passing that at $u=0$, where $\psi$, $\chi$ and either $R$ or $G$ or both are still exact, $\Xi\psi$ and $\Xi\psi$ converge to third order.
The results for the six combinations of initial Bondi gauge or initial affine gauge with the eR, eG and fe formulations are broadly similar.
In order to test the Raychaudhuri equation $E_{xx}=0$, we have also discretized
$E_{xx(1)}$ $:=$ $GV_{,x},$ (83)
$E_{xx(2)}$ $:=$ $-4\pi R(\psi_{,x}^{2}+\chi_{,x}^{2})$ (84)
using a different methods from elsewhere in the code. This means in the eR or eG formulations, where $G$ or $R$, respectively, are found by solving $E_{xx}=0$, we have a non-trivial test of the discretization of that process. In the fe formulation, we have a test of the discretization of the solution of $E_{xx}=0$ for either $R$ or $G$ in the initial data, combined with a test of how the subsequent free evolution error leads to increasing violation of the constraint $E_{xx}=0$. We find that $E_{xx(1)}-E_{xx(2)}=0$ converges to second order, except at the boundaries, where it is only first-order. This is because there we use second-order accurate but one-sided finite differences in evaluating the constraint, and therefore in self-convergence testing at the boundaries we compare a one-sided derivatives at the lower resolution with a centred derivative at the higher resolution.
For all variables, we have also taken the three differences of the numerical solutions in the three formulations, and again we find second-order convergence, with errors of the order of magnitude of the self-convergence errors in each formulation. In other words, all three formulations converge to the same continuum solution, as expected.
V.3 Evolving through the event horizon and singularity excision
We have further evolved the initial data that we already used for convergence testing through the horizon, obviously in eG and fe only. We lose pointwise convergence there, but the solutions in eG and fe remain qualitatively alike. On each time slice beyond the event horizon, $R$ increases from zero to a maximum that marks the apparent horizon, then decreases back to zero at some $x=x_{s}(u)$ which denotes the spacelike curvature singularity, with $dx_{s}/du<0$. We excise grid points with $x>x_{s}(u)$, which are no longer physical. The code remains stable and appears accurate up to within very few grid points of the singularity.
As $x\to x_{s}$, $R_{,x}$ and $G$ diverge. We therefore also plot against $u$ and $\lambda$. We find that $R$ then decreases essentially linearly with $\lambda$ everywhere inside the apparent horizon until we lose resolution near the singularity.
[…]
Clearly, our code is not yet very suitable for exploring aspects of charged black hole interiors such as Cauchy horizons and mass inflation, but that is not its purpose at the moment. Rather, we want to demonstrate that we can evolve into a region where outgoing null cones reconverge.
V.4 Type-II critical collapse of a spherical charged scalar field
As our emphasis in this paper is on developing methods for simulating type-II critical collapse on null cones beyond spherical symmetry, rather than on examining charged collapse in spherical symmetry, we examined only two one-parameter families of initial data, namely (82) with $x_{c}=1.5$, $d=0.5$, $\omega=2.0$ and ${\cal A}=0.05$, for either $R(0,x)=x$ or $G(0,x)=1$. We set the coupling constant to $q=1/\sqrt{4\pi}$. During the evolution, we use sdn gauge, given in (60), with the parameter $x_{0}$ adjusted so that $x=x_{0}$ is approximately the past light cone of the accumulation point of echos in near-critical solutions.
Numerically, we use either the eR or eG formulation. We have not tried the fe formulation in critical collapse, as we expect the slow drift away from the regularity condition $R_{,x}=G$ at the centre to affect the correctness of near-critical evolutions. As for numerical parameters, we use a uniform grid in $x$ with $x_{\text{max}}=5$ and $N_{x}=250$ grid points, and the timestep (79) with $C_{2}=0.1$ and $C_{3}=0.001$.
For the $G=1$ initial data, we chose $x_{0}=3.17$ in both the eR and eG formulation. We found $p_{*}\simeq 0.91$. The initial mass and charge at $p\simeq p_{*}$ are $M=0.151$ and $Q\simeq 0.021$.
For the $R=x$ initial data, we chose $x_{0}=3.107$ in the eR formulation and $x_{0}=3.103$ in the eG formulation. We found $p_{*}\simeq 1.01$. The initial mass and charge at $p\simeq p_{*}$ are $M=0.173$ and $Q\simeq 0.026$.
The results for the two families of initial data are qualitatively similar, as one would expect. The results for the eR and eG formulations, for the same initial data, are quantitatively very similar, as one would expect from the convergence tests. In the following, for conciseness we present results only for the $G=1$ family of initial data, evolved with the eG formulation.
As a proxy for the black hole (Hawking) mass and charge, which are strictly speaking defined at future null infinity, we evaluate $M$ and $Q$ on the first symmetry 2-sphere where the Hawking compactness takes the value $C(u,x)=C_{\text{thr}}$. (“First” here means the smallest such $u$.) In the results presented here in the eG formulation, we choose the obvious value $C_{\text{thr}}=1$, but this is not possible in the eR formulation. We discuss the role of the choice of $C_{\text{thr}}$ in Sec. V.6.
Figure 1: Maximum curvature fine-structure for the $G=1$ family of initial data, in the eG formulation: we plot $1/2\ln|\max T|-\gamma\ln(1-p/p_{*})+A$ against $\ln(1-p/p_{*})$. We use the approximate theoretical value $\gamma=0.374$, and the family-dependent offset A=0.23A=-0.23 has been fitted by eye so that the average of the fine-structure is approximately zero at high fine-tuning. The range of $\ln(1-p/p_{*})$ on the horizontal axes has been chosen as in Fig. 1 of Paper II for comparison. (As in Paper II, we have suppressed axis labels to make figures larger.) The thin dotted line has been fitted by eye to the first two peaks from the right and has slope $\delta\gamma=0.012$, meaning that the curvature critical exponent is slightly larger at low fine-tuning.
Figure 2: Black hole mass fine-structure for the $G=1$ family of initial data, in the eG formulation, with $C_{\text{thr}}=1$: we plot $\ln M-\gamma\ln(p/p_{*}-1)+A$ against $\ln(p/p_{*}-1)$. $\gamma$ and $A$ and the horizontal range are as for the curvature scaling plot in Fig. 1. Compare also Fig. 2 of Paper II. The thin dotted line has been fitted by eye to the first two peaks from the right. It has slope $\delta\gamma=0.010$, meaning that the mass critical exponent is slightly larger at low fine-tuning, in line with that for that of the curvature to within a plausible fitting error.
Figure 3: Black hole charge fine-structure for the $G=1$ family of initial data, in the eG formulation, with $C_{\text{thr}}=1$: we plot $\ln Q-\mu\ln(p/p_{*}-1)+A$ against $\ln(p/p_{*}-1)$. We have used the approximate expected critical exponent $\mu=0.883$, and AA as in the curvature scaling plot in Fig. 1. The thin dotted line has been fitted by eye to the first two peaks from right. It has slope $\delta\mu=-0.059$, meaning that the charge critical exponent is smaller at low fine-tuning. It is plausible from this figure that at high fine-tuning the expected value of $\mu$ is achieved. The fine-structure has approximately the same period as for the curvature and mass, as one expects.
Figure 4: Maximum curvature fine-structure for the $G=1$ family of real initial data, in the eG formulation. We set A=0.58A=-0.58, otherwise as in Fig. 1.
Figure 5: Black hole mass fine-structure for the $G=1$ family of real initial data, in the eG formulation, with $C_{\text{thr}}=1$. We set A=0.58A=-0.58, otherwise as in Fig. 2.
The fine-structure of critical scaling on the dispersal side of the threshold of collapse is shown in Fig. 1 for the curvature invariant $T$. On the collapse side, Figs. 2 and 3 show the fine-structure for $M$ and $Q$ of the first (approximate) MOTS. What we actually plot is $1/2\ln T-\gamma\ln(1-p/p_{*})+A$ in Fig. 1, $\ln M-\gamma\ln(1-p/p_{*})+A$ in Fig. 2, and $\ln Q-\mu\ln(1-p/p_{*})+A$ in Fig. 3. The family-dependent offset $A$ has been fitted so the average of $1/2\ln T-\gamma\ln(1-p/p_{*})+A$ is roughly zero at high fine-tuning. $\gamma$ and $\mu$ have been set to their theoretically predicted approximate values $\gamma=0.374$ and $\mu=0.883$ GundlachMartin96 .
V.5 The effect of charge on the critical scaling
At high fine-tuning, the fine-structures are periodic, thus confirming the values of $\gamma$ and $\mu$ from perturbation theory. The fine structure for $T$ has the same amplitude of about $0.15$ and flat top as in Fig. 1 of Paper II, which showed $T$ for a one-parameter family of real, spherically symmetric initial data. At low fine-tuning, the curve has the same periodic structure but rises slightly with $\ln(1-p/p_{*})$. The same observation applies to the periodic fine structure of $M$ at high fine-tuning. It has an amplitude $\simeq 0.03$. From Fig. 3, the fine structure in $\ln Q$ evaluated also appears to be continuous, with amplitude $\simeq 0.15$.
The value of $\mu$ predicted in GundlachMartin96 was verified in collapse simulations in HodPiran97 and Petryk05 , and is verified here again by us. Figures in these last two papers showed a power law with a wiggle, but did not show the wiggle clearly. Hence our Fig. 3 is the first plot of the fine structure of the charge scaling law for any physical system, here the spherical charged scalar field.
At low fine-tuning, we see deviations from the expected scaling laws. Heuristically, these can be described as modifications of the critical exponents $\gamma$ and $\mu$ and the echoing period $\Delta$ (as we did for the critical exponents for the mass and angular momentum in GunBau18 ). To quantify these modifications, we have fitted the thin dotted lines in Figs. 1, 2 and Fig. 3 by eye. These have slopes of $\delta\gamma=0.012$, $\delta\gamma=0.010$, and $\delta\mu=-0.059$, respectively. Note the values of $\delta\gamma$ on the dispersion and collapse sides are roughly the same, as one would expect, but the fitting by eye is somewhat arbitrary. As in GunBau18 , the small change in critical exponent is a convenient quantitative way of describing the fine structure at low fine-tuning, but we do not have a quantitative explanation for it.
For comparison, we have also fine-tuned a family of uncharged $G=1$ initial data, differing from our charged data only by the choice $\omega=0$. We also set ${\cal A}=0.075$ in order to make $p_{*}\simeq 1$. For the eG formulation, we chose $x_{0}=3.525$ and found $p_{*}\simeq 0.95$. For the eR formulation, we chose $x_{0}=3.530$ and found $p_{*}\simeq 0.94$. The fine-structure for $T$ and $M$ (in the eG formulation, with $C_{\text{thr}}=1$) is given in Figs. 4 and 5. Here we can read off more accurately that the amplitude of the fine structure for $M$ is $\simeq 0.032$. As already noted in Paper II, this differs significantly from the values for the amplitude given in the literature. It is remarkable how periodic both fine-structures are already at low fine-tuning, in contrast to the charged case, where $|Q|/M$ has to be quite small before the critical exponents agree with their values calculated in perturbation theory in GundlachMartin96 . That calculation considered charge as a linear perturbation on an uncharged background critical solution (the Choptuik solution). From Figs. 1, 2 and 3, this seems to be accurate for $\ln|1-p/p_{*}|\lesssim-20$. This corresponds to $|Q|/M\lesssim 10^{-4}$.
Here we present results only for a single one-parameter family of charged initial data, which tells us nothing about universality. Both the mass and charge critical exponents appear to agree with their predicted values, however, which is indirect evidence of their universality. We also find that the fine-structure of the mass scaling law looks very similar in the charged and uncharged families. By contrast, we have found initial evidence that the fine-structure of the charge scaling law may not be universal. We will return to this elsewhere.
V.6 The role of the collapse criterion
Figure 6: The fine structure of $M$, for the $G=1$ family of initial data, in the eG formulation, but now for different values of $C_{\text{thr}}$. From bottom to top, the curves correspond to $C_{\text{thr}}=0.8$ $0.9$, $0.95$, $0.99$, $0.999$ and $1$, as in Fig. 7. The $C_{\text{thr}}=0.999$ and 11 curves are almost indistinguishable, but there is a tiny jump once per period in the $C_{\text{thr}}=0.999$ curve. The $C_{\text{thr}}=1$ curve is the same as in Fig. 2.
Figure 7: Location $x=x_{C}$ of the symmetry sphere $(u_{C},x_{C})$ where we evaluate $M$ and $Q$, for the $G=1$ family of initial data, in the eG formulation, but now for different values of $C_{\text{thr}}$. From bottom to top, the curves correspond to $C_{\text{thr}}=0.8$, $0.9$, $0.95$, $0.99$, $0.999$ and $1$.
In the eG and fe formulations we can choose $C_{\text{thr}}=1$, but not in the eR formulation. In Paper II, for example, where we used the eR formulation, we used $C_{\text{thr}}=0.99$ in spherical symmetry and for mildly non-spherical initial data (and $C_{\text{thr}}=0.8$ for the most-nonspherical initial data), and found that the mass fine-structure was discontinuous, including in spherical symmetry.
To demonstrate the effect of the value of $C_{\text{thr}}$ on the measured $M$ and $Q$, Fig. 6 shows the mass fine-structure obtain with different values of $C_{\text{thr}}$. From bottom to top, the six curves correspond to $C_{\text{thr}}=0.8$, $0.9$, $0.95$, $0.99$, $0.999$ and $1$ of $C$. The curves for $C_{\text{thr}}<1$ are discontinuous once per period, but the curve for $C_{\text{thr}}=1$ is continuous. (The jump in the curve for $C_{\text{thr}}=0.999$ is just about visible.) Similar jumps are seen in the fine-structure of $Q$.
Fig. 7 shows the $x$-location of the first symmetry 2-sphere $(u_{C},x_{C})$ where $C(u_{C},x_{C})=C_{\text{thr}}$, for the same values of $C_{\text{thr}}$. For each $C_{\text{thr}}$, this location is discontinuous at the same values of $\ln(p/p_{*}-1)$ where $\ln M$ and $\ln Q$ also jump. The jump in $x_{C}$ is still large for $C_{\text{thr}}=0.999$. This suggests that, with $C_{\text{thr}}<1$, the jumps in $M$ and $Q$ at periodic values of $\ln(p/p_{*}-1)$ are caused by the jumps in $x_{C}$ at the same values.
The exception is the location $x_{C}$ of the first occurrence of $C=1$: it is a very noisy function of $\ln(p/p_{*}-1)$, but with no obvious jumps. The noisiness is explained, in hindsight, by the difficulty of locating the first occurrence of $C=1$: in practice, we find the $x$-location where $C$ goes through 0 at the first timestep where its maximum is larger than 1. Note that $M$ and $Q$ read off at this noisy location are not visibly noisy in Figs. 2 and 3: we believe this is explained by the fact that $M$ and $Q$, as functions of $x$ along the curve $C=1$ quickly asymptote to $M_{\text{BH}}$ and $Q_{\text{BH}}$, and so the values of $M$ and $Q$ read off at $x=x_{C}$ depend only weakly on $x_{C}$.
We can also understand theoretically why the mass fine-structure in spherical symmetry is continuous for $C_{\text{thr}}=1$ but discontinous for $C_{\text{thr}}<1$. We expect the function $C(u,x;p)$ to be continuous in all its arguments. Nevertheless, the set of points $C(u,x;p)=C_{\text{thr}}$ changes discontinously in $p$ whenever a local maximum of $C(u,x)$ rises above $C_{\text{thr}}$, and a new contour line forms. Therefore the location of the first 2-sphere where $C(u,x;p)=C_{\text{thr}}$ also depends discontinuously on $p$. $C_{\text{thr}}=1$ is an exception because the only contour line of $C(u,x;p)=1$ is the apparent horizon, dividing spacetime into a trapped and an untrapped region. But this exception does not apply beyond spherical symmetry, where a 2-sphere with $C=1$ is not in general marginally trapped.
Based on our numerical experiments with different values of $C_{\text{thr}}$, and this theoretical understanding, we now believe that the discontinuity in the mass fine-structure observed in Paper II was an artifact of the choice $C_{\text{thr}}=0.99$, at least in spherical symmetry. Beyond spherical symmetry, we are not sure. We will return to this question elsewhere.
As a practical discovery, we find that plotting the equivalent of Fig. 7 just for the data points obtained in bisection allows us to quickly find the appropriate value of $x_{0}$: if these curves drop to $x=0$ with increased fine-tuning, our choice of $x_{0}$ is too large (and we lose resolution in the self-similar region). If they rise to $x=x_{\text{max}}$ then $x_{0}$ is too small (and the first occurence of the specified threshold value of $C$ actually occurs at larger $x$, outside our numerical domain).
VI Tests on a null rectangle
VI.1 Previous work
We now test evolutions with null data given on an intersecting outgoing null cone $u=u_{0}$, for $v_{0}\leq v\leq v_{1}$ (the “right boundary”), and an ingoing null cone $v=v_{0}$ , for $u_{0}\leq u\leq u_{1}$ (the “left boundary”). These data determine the solution on the “null rectangle” $u_{0}\leq u\leq u_{1}$, $v_{0}\leq v\leq v_{1}$.
Simulations on null rectangles have been used in Gnedin93 ; BradySmith95 ; MurReaTan13 to study black hole interiors in the simpler case of the spherical Einstein-Maxwell-real scalar field system.
By contrast, in GellesPretorius25 , the spherical Maxwell-charged scalar system is evolved on a fixed RN spacetime, in order to investigate event horizon instabilities. That work uses double-null coordinates, compactified to both past and future null infinity (but finite across the past and future event horizon) as in MurReaTan13 . It uses the electromagnetic gauge $\nabla_{a}(R^{-2}A^{a})=0$, or $A_{u,v}+A_{v,u}=0$ in double-null coordinates. (This is in contrast to our electromagnetic gauge $A_{x}=0$, which in double-null coordinates becomes $A_{v}=0$.)
We are aware of only one previous numerical treatment of the full Einstein-charged scalar system in spherical symmetry: in BaakeRinne16 , the evolution is on hyperboloidal spacelike slices that on the left end are truncated with an excision boundary inside the event horizon of a RN black hole, and at their right end extend to future null infinity.
We believe the present paper gives the first algorithm for simulating the spherical Einstein-Maxwell-charged scalar system in null coordinates, allowing for simulations both on a null rectangle or with a regular centre.
VI.2 Physical setup
[…]
Our algorithm also allows us to fix the right spacetime gauge as $R_{,x}=R_{,x*}$ (so that $x$ is a linear funtion of $R$), but for brevity we do not present results here. When there is no scalar field on the right boundary, Bondi and affine right gauge are equivalent anyway. Morever, in our discretisation they are equivalent without discretisation error.
We can continue the spacetime gauge into the numerical domain with any choice of $B(u,x)$ (as long as $B=0$ on the left boundary), but again for brevity we only present double-null gauge $B(u,x)=0$, so that $x$ is an ingoing null coordinate everywhere. We use the timestep $\Delta u_{0}=C_{0}\Delta x$, with $C=0.1$.
We shall restrict to scenarios where $C_{*}<1$, so that the corner 2-sphere is untrapped. Then $R_{,x*}>0$ gives us $R_{,u*}<0$, and the ingoing Raychaudhuri equation (49) gives us $R_{,u}<0$ on the entire left boundary. We also restrict to the case where the right boundary is untrapped, so that $R_{,x}>0$ there.
VI.3 Prompt collapse
Figure 8: Spacetime diagram of the prompt collapse scenario, showing the regular centre and numerical domain in bold, the singularity in bold-dashed, and the event horizon as a thin line.
Figure 9: Spacetime diagram of collapse from data on a null cone with regular centre, with the lines as in Fig. 8. For suitable data, this could be the same spacetime as in Fig. 8.
Consider a null rectangle with $R_{*}>0$, $R>0$ on the left boundary, $M_{*}=Q_{*}=0$, zero scalar field on the left boundary, and no trapped surface on the right boundary. Then the triangular region bounded by the regular centre $R=0$, $u=u_{0}$ and $v=v_{0}$ is a piece of Minkowski spacetime, and $u=u_{0}$ can be extended to the left to $R=0$ and to the right to $R=\infty$ as a lightcone with a regular vertex and no trapped surface.
If the data on $u=0$ are sufficiently strong that a MOTS first forms at $(v_{1},u_{\text{FMOTS}})$ for some $u_{0}<u_{\text{FMOTS}}<u_{1}$, then there must be an event horizon at some $u=u_{\text{EH}}$ with $u_{0}<u_{\text{EH}}\leq u_{\text{FMOTS}}$. We have in effect simulated gravitational collapse from initial data on a null cone with regular centre, while actually solving the Einstein equations only on a null rectangle, and thus avoiding the complications of a regular centre. The null rectangle may then be truncated by a curvature singularity “at the top”. Figs. 8 and 9 show gravitational collapse in the null rectangle and regular centre settings.
If, on the other hand, there is no trapped surface in the null rectangle, formation of a black hole from the regular data on $u=u_{0}$ is neither proved nor disproved. In particular, we do not expect to find the black hole threshold in a one-parameter family of initial data on $u=u_{0}$ in this setup: $u_{\text{FMOTS}}$ will become larger than $u_{1}$ well above the collapse threshold. Another way of seeing this is to note that the critical solution in type-II critical collapse is non-vacuum both inside and outside the past null cone of its singularity.
As an example of prompt collapse, we set ${\cal M}_{*}=Q_{*}=0$, $R_{*}=2$, and fix the corner gauge as $G_{*}=R_{,x*}=1$. With no scalar field on the left boundary, $R_{,u}=-1/2$ there, so the left boundary intersects $R=0$ at $u=4$. On the right boundary we set $\psi$ to be a Gaussian with centre at $x_{c}=0.5$, width $d=0.1$, frequency $\omega=50$ and amplitude $9.5\cdot 10^{-3}$. We set $q=50$. From now on, without loss of generality, we let $u_{0}=0$ and $v_{0}=v$, and we identify $v=0$ with $x=0$. The event horizon forms just after $u=3.3$. In the eG and fe formulations we evolve to $u=3.9$, just before singularity excision starts, and in the eR formulation at $u=3.3$, just before the event horizon forms.
We evolve with $N_{x}=100\cdot 2^{n}$ grid points, with n=04n=0...4. In all three formulations we find perfect second-order pointwise self-convergence of $R$ and $\ln G$ already at the lowest resolution, without any of the complications caused by using expansions at a regular centre. Perfect second-order self-convergence persists through the event horizon and until excision starts, and even during excision breaks down only near the excision boundary. The pairwise differences between the three formulations similarly show perfect second-order pointwise convergence.
VI.4 Schwarzschild and extremal Reissner-Nordström
In our gauge, the vacuum Schwarzschild solution is already a non-trivial test, but with an exact solution available, see Appendix A. With the choice $G_{*}=R_{,x*}=1$, affine and Bondi right gauge coincide, with $G=R_{,x}=1$ on the entire right boundary. We set $M_{*}=0.5$, so that the horizon radius is $r_{+}=1$, and $R_{*}=2$, so that $C_{*}=1/2$ and $R_{,u}=R_{,u*}=-1/4$ on the entire left boundary. We set $x_{\text{max}}=1$. In the eR formulation we stop the evolution at $u=3.9$, just before the event horizon $R=1$ is crossed at $u=4$. In the eG and fe formulations, singularity excision starts at $u\simeq 4.3$, and we stop the evolution at $u=6$, when all grid points at $x\gtrsim 0.1$ are excised. We again evolve with $N_{x}=100\cdot 2^{n}$, with n=04n=0...4. We see perfect self-convergence and convergence between resolutions to the end of the simulations, that is close to the horizon in eR and well into the excision regime in eG and fe. We also see perfect convergence of all formulations against the exact solution (we have coded up the exact solution only outside the horizon).
We continue to extremal RN. We set ${\cal M}_{*}=Q_{*}=1$, so that the horizon radius is again $r_{+}=1$, and we again set $R_{*}=2$. $M$ is now not constant, and in particular at the corner $M_{*}=3/4$, so that $C_{*}=3/8$. We again fix the gauge as $G_{*}=R_{,x*}=1$. This gives us $R_{,u}=-1/8$ on the left boundary, so that the event horizon is crossed at $u=8$, and the left boundary runs into the singularity at $u=16$.
On the horizon $R_{,x}=0$ but, in contrast to the subextremal case, inside the black hole $R_{,x}>0$ again, and the singularity $R=0$ is now timelike. This means that our time slices of constant $u$ run into the singularity at their left edge first. Singularity excision would not make sense, as boundary conditions would have to be imposed at the timelike singularity.
We evolve with $N_{x}=100\cdot 2^{n}$, with n=06n=0...6, up to $u=7.9$ in the eR formulation, and up to $u=14$ in the eG and fe formulations. We see perfect self-convergence and convergence between formulations in the highest resolutions, but convergence in the lower resolutions begins to break down between $u=13$ and $u=14$ due to the appearance of sharp gradients in all variables near the left boundary.
VI.5 Perturbed extremal Reissner-Nordström
To simulate extremal RN perturbed by charged matter entering through the right boundary, we set again ${\cal M}_{*}=Q_{*}=1$, $R_{*}=2$, $G_{*}=R_{,x*}=1$, $x_{\text{max}}=1$, and we evolve to $u=14$, this time only in the eG and fe formulations as we want to cross the horizon.
We set $\phi=0$ on the left boundary, but on the right boundary we set $\phi$ to a complex Gaussian of the form (82) with centre at $x_{c}=0.4$, width $d=0.1$, complex phase frequency $\omega=30$ and amplitude $5\cdot 10^{-4}$. We set $q=100$. Then at $u=0$ the augmented mass rises from ${\cal M}_{*}=1$ at $x=0$ to ${\cal M}\simeq 1.0032$ at $x_{\text{max}}=1$ (the outer boundary), and the charge from $Q_{*}=1$ to $Q\simeq 1.0068$. One can interpret this as trying to overcharge a pre-existing static eRN black hole with a burst of scalar field radiation coming in from past null infinity at finite advanced time.
Not unexpectedly, this particular attempt fails. Indeed for all $u>1.7$, already well before the putative horizon is crossed, $Q\leq{\cal M}\leq 1$ everywhere on the domain $0\leq x\leq 1$ (with equality only at the left boundary, where the solution is still electrovacuum extremal RN). The event horizon is crossed at $u\simeq 9$ and has radius $R\simeq 0.9$ (and hence Hawking mass $M\simeq 0.45$), augmented mass ${\cal M}\simeq 0.996$ and charge $Q\simeq 0.991$, so the final black hole is slightly subcritical.
However, while $R$ on our numerical domain $0\leq x\leq 1$ levels off in $x$ at $u\simeq 9$, it returns to being an increasing function of $x$ afterwards, at least up to $u=14$, when we stop the simulation. Although this behaviour is characteristic of an extremal black hole (which has no trapped surfaces in the interior), we know in this case that the black hole is clearly subcritical. Hence the fact that we do not see $R_{,x}<0$ or the spacelike singularity $R=0$ on our numerical domain must be due to its limited size.
We evolve with $N_{x}=100\cdot 2^{n}$, now with n=06n=0...6. We need the higher resolutions to see clean self-convergence and convergence between resolutions. The simulation is more demanding than that of unperturbed eRN because $\phi$ becomes more rapidly oscillating in $x$: in the initial data, $\phi$ has a about two complex oscillations within a Gaussian envelope, while at $u=14$ there are about 30 complex oscillations in roughly the same envelope.
[…]
Once again we see good convergence within formulations and between formulations at the highest resolution. At $u=14$, the complex scalar field is rapidly oscillating in $x$, with about 20 phase oscillations between $x=0$ and $x=0.1$, with an exponentially decaying envelope, and this is only just resolved at the two highest resolutions.
VII Conclusions
This series of papers investigates the usefulness of “Bondi-like” null coordinates, in which surfaces of constant retarded time $u$ are null cones and the coordinate lines of constant $(u,\theta,\varphi)$ are their generators. The remaining gauge freedom is how points on each generator are labelled by the radial coordinate $x$.
We have identified three generic formulations: the metric coefficient $R$ is evolved and the metric coefficient $G$ is found by solving the (outgoing) Raychaudhuri equation, or the other way around, or both $R$ and $G$ are evolved freely (and the Raychaudhuri equation is imposed as a constraint only at $u=0$). In these three, any labelling of points on a null generator by the radial coordinate $x$ is possible. Dynamically, this radial gauge choice is controlled by freely choosing $B$. There are also two particular formulations each tied to a specific gauge: the Bondi formulation where $R=x$, and the affine formulation where $G=1$.
To our knowledge, there is no clear statement in the literature that any of the three general formulations can be combined with any choice of radial gauge. This observation is the main conceptual novelty of the present paper.
To present the five different formulations pedagogically, and with a focus on regularity at the central worldline $R=0$, in this paper we have restricted attention to spherical symmetry. We will explore the eG formulation in axisymmetry in a future paper, but we notice already that the generalisation beyond spherical symmetry and/or to higher dimensions is conceptually straightforward.
As a nontrivial test-bed in spherical symmetry, we have implemented the spherical Einstein-Maxwell-charged scalar system, allowing for initial data either on a single outgoing null cone $u=0$ emerging from a regular centre $R=0$, or on two intersecting null cones $u=u_{0}$ and $v=v_{0}$.
In both setups we have shown second-order self-convergence in all three formulations and second-order convergence between all three formulations. With a regular centre, the convergence is pointwise except near the centre, where it still holds in any norm, such as the root-mean-square or maximum norm. The fact that even the fe formulation converges without any tweaks at the centre came as a pleasant surprise. With the left boundary an ingoing null cone, the convergence is perfectly pointwise everywhere.
As a physics testbed of collapse from regular data on an outgoing null cone, we have investigated critical collapse of a charged scalar field. This had previously been investigated theoretically in GundlachMartin96 , and numerically in HodPiran97 ; Petryk05 , but here we show the first accurate plots of the universal fine structure in the mass and charge scaling laws. Surprisingly, we found that being able to evaluate the mass and charge on a marginally trapped surface makes a qualititative difference to our results, compared to evaluating them on an almost-trapped surface, even one with Hawking compactness as large as $C=0.999$. We also found initial evidence that the fine-structure of the charge scaling law may not be universal. We will return to this elsewhere.
However, the main motivation for introducing the two formulations in which $G$ is evolved is that in them we can evolve through event and apparent horizons. In Paper II, both our formulation (we evolved $R$) and our choice of gauge (“local shifted Bondi gauge”) did not allow us to do this in axisymmetric scalar collapse. This in turn meant that we could not bisect arbitrarily close to the collapse threshold in highly-nonspherical one-parameter families of initial data. We plan to apply our eG formulation with a suitable radiable gauge to this problem.
As a physics testbed of the setup with initial data on two intersecting null cones, we have simulated gravitational collapse with a Minkowski interior, the exact Schwarzschild and extreme Reissner-Nordström solutions (which are non-trivial in our gauge), and extreme Reissner-Nordström perturbed by a complex scalar field on either $u=u_{0}$ or $v=v_{0}$.
We plan a study of the solution space of the spherical Einstein-Maxwell-charged scalar system, and in particular an investigation of the “extremal critical collapse conjecture for the spherical Einstein-charged scalar field system KehleUnger24 , see also KehleUnger22 . We anticipate that both types of numerical domain will be required for this investigation.
Appendix A The Reissner-Nordström metric in null coordinates
This Appendix presents the RN solution in our variables and gauge choice. In Schwarzschild coordinates, where $R=r$ is a coordinate, the RN metric is
$ds^{2}=-f\,dt^{2}+f^{-1}\,dr^{2}+r^{2}\,d\Omega^{2},$ (85)
where
$f(r):=1-{2{\cal M}_{0}\over r}+{Q_{0}^{2}\over r^{2}},$ (86)
and the real constants ${\cal M}_{0}>0$ and $Q_{0}$ are the usual mass and charge parameters of the RN metric, normalised such that $|Q_{0}|={\cal M}_{0}$ represents an extremal black hole. The event and Cauchy horizon are at
$r_{\pm}={\cal M}\pm\sqrt{{\cal M}^{2}-Q^{2}}.$ (87)
By definition the Hawking mass obeys $r=2M$ on the event horizon. $M=r_{+}/2$ at the event horizon is also the irreducible mass of the black hole. In the special case of an extremal black hole, $r_{+}=r_{-}=2M={\cal M}=Q$ on the event horizon.
The tortoise radius $r_{*}$ is defined, up to addition of a constant, by
${dr_{*}\over dr}:={1\over f(r)}.$ (88)
With the outgoing and ingoing Eddington-Finkelstein coordinates defined as
$\bar{u}:=t-r_{*},\quad\bar{v}:=t+r_{*},$ (89)
the metric in the exterior becomes
$ds^{2}$ $=$ $-f\,d\bar{u}\,d\bar{v}+r^{2}\,d\Omega^{2}$ (90)
$=$ $-f\,d\bar{u}^{2}-2\,d\bar{u}\,dr+r^{2}\,d\Omega^{2}$ (91)
$=$ $-f\,d\bar{v}^{2}+2\,d\bar{v}\,dr+r^{2}\,d\Omega^{2}.$ (92)
(Both $\bar{v}$ and $\bar{u}$ increase to the future. In the interior, we define $\bar{u}:=r_{*}-t$ instead to achieve this.)
Writing (91) in our notation, and choosing our coordinate $x$ to coincide with $r$, the RN metric in Bondi gauge has the metric coefficients
$R$ $=$ $x,$ (93)
$G$ $=$ $1,$ (94)
$B$ $=$ ${1\over 2}\left(1-{2{\cal M}_{0}\over x}+{Q_{0}^{2}\over x^{2}}\right).$ (95)
[…]
$R$ $=$ $r\left({\bar{v}-\bar{u}}\over 2\right),$ (96)
$\bar{G}$ $=$ ${1\over 2}f(R),$ (97)
$B$ $=$ $0,$ (98)
where $r(r_{*})$ is defined as the function inverse of $r_{*}(r)$, and $\bar{G}$ denotes the value of $G$ in the particular double-null coordinates $\bar{u}$, $\bar{v}$.
If, for given constants $\bar{u}_{0}$, $\bar{v}_{0}$, $u_{0}$, $v_{0}$, we now define the alternative double-null coordinates
$u(\bar{u})$ $:=$ $u_{0}-a\left(r\left({\bar{v}_{0}-\bar{u}\over 2}\right)-r_{0}\right),$ (99)
$v(\bar{v})$ $:=$ $v_{0}+b\left(r\left({\bar{v}-\bar{u}_{0}\over 2}\right)-r_{0}\right),$ (100)
where
$r_{0}:=r\left({\bar{v}_{0}-\bar{u}_{0}\over 2}\right),$ (101)
we have $u(\bar{u}_{0})=u_{0}$, $v(\bar{v}_{0})=v_{0}$, $R_{,u}=-1/a$ along $v=v_{0}$ and $R_{,v}=1/b$ along $u=u_{0}$. Inverting, we obtain
$\bar{u}(u)$ $=$ $\bar{v}_{0}-2r_{*}\left(r_{0}+{u-u_{0}\over-a}\right),$ (102)
$\bar{v}(v)$ $=$ $\bar{u}_{0}+2r_{*}\left(r_{0}+{v-v_{0}\over b}\right),$ (103)
$R(u,v)$ $=$ $r\Bigl[r_{*}\left(r_{0}+{v-v_{0}\over b}\right)$ (104)
$+r_{*}\left(r_{0}+{u-u_{0}\over-a}\right)-r_{*}(r_{0})\Bigr].$
We have
${du\over d\bar{u}}$ $=$ $a\,\bar{G}(\bar{u},\bar{v}_{0}),$ (105)
${dv\over d\bar{v}}$ $=$ $b\,\bar{G}(\bar{u}_{0},\bar{v}),$ (106)
and therefore
$G(u,v)={\bar{G}(\bar{v},\bar{u})\over ab\,\bar{G}(\bar{v}_{0},\bar{u})\,\bar{G}(\bar{v},\bar{u}_{0})},$ (107)
and hence $G(u_{0},v)=G(u,v_{0})=G_{*}$. This is, in implicit form, the exterior metric of RN in the coordinates we use in our null-rectangle tests, with $(R_{,v})_{*}=1/b$, $R_{,u*}=-1/a$, and
$G_{*}={1\over ab\,\bar{G}(\bar{u}_{0},\bar{v}_{0})}.$ (108)
Appendix B Regular centre
The Minkowski metric in Bondi coordinates, with $U$ normalised to be proper time at the centre, is
$ds^{2}=-dU^{2}-2dU\,dR+R^{2}\,d\Omega^{2}.$ (109)
Now let $U=U(u)$ and $R=R(u,x)$. Our metric coefficients become
$G$ $=$ $U^{\prime}R_{,x},$ (110)
$B$ $=$ ${U^{\prime 2}+2U^{\prime}R_{,u}\over 2U^{\prime}R_{,x}},$ (111)
$\Xi R$ $=$ $-{U^{\prime}\over 2},$ (112)
${\cal H}$ $=$ $-{U^{\prime\prime}\over U^{\prime}}.$ (113)
We impose $R(u,0)=0$, thus fixing the regular centre at $x=0$, and we make $u$ proper time at the centre by setting $U=u$. This gives us
$G(u,0)$ $=$ $R_{,x}(u,0),$ (114)
$B(u,0)$ $=$ ${1\over 2R_{,x}(u,0)},$ (115)
$\Xi R(u,0)$ $=$ $-{1\over 2},$ (116)
${\cal H}(u,0)$ $=$ $0.$ (117)
From elementary flatness, all these conditions at the centre must also hold in an arbitary curved spherical spacetime with a regular centre $R=x=0$ and $u$ proper time at the centre. Beyond spherical symmetry, in spherical coordinates, they will also hold at the regular “central” worldline $R=x=0$ as long as it is geodesic, see Appendix III.E.2 of Paper I for this subtlety.
Appendix C Expansions at the centre
For finding $G$ by integration, given $R$, we need
$\ln\left({G\over R_{,x}}\right)$ $=$ $2\pi(\psi_{(1)}^{2}+\chi_{(1)}^{2})x^{2}$ (118)
$+{4\pi\over 3}\Bigl[4(\psi_{(1)}\psi_{(2)}+\chi_{(1)}\chi_{(2)})$
$-{R_{(2)}\over R_{(1)}}(\psi_{(1)}^{2}+\psi_{(2)}^{2})\Bigr]x^{3}+O(x^{4}).$
Equivalently, for finding $R$ by solving a system of two first-order ODEs, given $G$, we require
$R$ $=$ $G_{(0)}x+{G_{(1)}\over 2}x^{2}$
$+{1\over 3}\left(G_{(2)}-2G_{(0)}\pi(\psi_{(1)}^{2}+\chi_{(1)}^{2})\right)x^{3}+O(x^{4}),$
$V$ $=$ $1-2\pi(\psi_{(1)}^{2}+\chi_{(1)}^{2})x^{2}$ (120)
$+{2\pi\over 3G_{(0)}}\Bigl[G_{(1)}(\psi_{(1)}^{2}+\chi_{(1)}^{2})$
$-8G_{(0)}(\psi_{(1)}\psi_{(2)}+\chi_{(1)}\chi_{(2)})\Bigr]x^{3}+O(x^{4}).$
Note that, given the definition of $V:=R_{,x}/G$, to be consistent we should truncate the expansion of $V$ at one order lower in $x$ than that of $R$, but here we have written out all expansion coefficients that we know in full given a quadratic fit to the evolved quantities. We note in passing that the reverse of (C) is
$G$ $=$ $R_{(1)}+2R_{(2)}x+\left(3R_{3}+2\pi R_{(1)}(\psi_{(1)}^{2}+\chi_{(1)}^{2})\right)x^{2}$ (121)
$+O(x^{3}).$
We next have
${Q\over{4\pi}q}$ $=$ ${R_{(1)}^{2}\over 3}(\chi_{(1)}\psi_{(0)}-\psi_{(1)}\chi_{(0)})x^{3}$ (122)
$+{R_{(1)}\over 2}\Bigl[R_{(2)}(\psi_{(0)}\chi_{(1)}-\chi_{(0)}\psi_{(1)})$
$+R_{(1)}(\psi_{(0)}\chi_{(2)}-\chi_{(0)}\psi_{(2)})\Bigr]x^{4}+O(x^{5}),$
${A\over{4\pi}q}$ $=$ ${R_{(1)}\over 6}(\chi_{(1)}\psi_{(0)}-\psi_{(1)}\chi_{(0)})x^{2}$ (123)
$+{1\over 6}\Bigl[R_{(2)}(\psi_{(0)}\chi_{(1)}-\chi_{(0)}\psi_{(1)})$
$+R_{(1)}(\psi_{(0)}\chi_{(2)}-\chi_{(0)}\psi_{(2)})\Bigr]x^{3}+O(x^{4}).$
Note that $R_{(3)}$ (or $G_{(2)}$) does not appear to this order.
The expansion of $\Xi R$ is
$\Xi R$ $=$ $-{1\over 2}-{\pi\over 3}(\psi_{(1)}^{2}+\chi_{(1)}^{2})x^{2}$ (124)
$-{2\pi\over 3}(\psi_{(1)}\psi_{(2)}+\chi_{(1)}\chi_{(2)})x^{3}+O(x^{4}),$
where we have already used the expansion of either $R$ or $G$ above, but the expansion of $R\Xi R$ takes the simpler form
$R\,\Xi R=-{G_{(0)}\over 2}x-{G_{(1)}\over 4}x^{2}-{G_{(2)}\over 6}x^{3}+O(x^{4}),$ (125)
The two are of course equivalent for analytic functions, but we use the latter in the code.
As one would expect, we can expand $R\hat{\Xi}\psi$ to one order higher than $\hat{\Xi}\psi$, and similarly for $\hat{\Xi}\chi$. The latter expansion is simpler, and is
$R\,\hat{\Xi}\psi$ $=$ ${\psi_{(1)}\over 2}x+{\psi_{(2)}\over 2}x^{2}+O(x^{3}),$ (126)
$R\,\hat{\Xi}\chi$ $=$ ${\chi_{(1)}\over 2}x+{\chi_{(2)}\over 2}x^{2}+O(x^{3}).$ (127)
Again none of the expansion coefficients for $R$ (or $G$) apear to this order.
Finally, we have
${\cal H}$ $=$ ${8\pi\over 3R_{(1)}}(\psi_{(1)}^{2}+\chi_{(1)}^{2})x$ (128)
$-{4\pi\over 3}\Biggl[{R_{(2)}\over R_{(1)}^{2}}(\psi_{(1)}^{2}+\chi_{(1)}^{2})$
$-{3\over R_{(1)}}(\psi_{(1)}\psi_{(2)}+\chi_{(1)}\chi_{(2)})\Biggr]x^{2}+O(x^{3}).$
Appendix D Radial null geodesics peeling off the horizon
We consider radial null geodesics in the RN metric. From (88) and (89) we have
$r_{,\bar{v}}(\bar{u},\bar{v})={1\over 2}f(r).$ (129)
Linearising about the event horizon $r=r_{+}$, we have
$r_{,\bar{v}}\simeq\kappa(r-r_{+}),\quad\kappa:={1\over 2}f^{\prime}(r_{+}),$ (130)
where $\kappa$ is the surface gravity. In the subextremal case $\kappa>0$ the surfaces of constant $u$, and hence the outgoing radial null geodesics that generate them, peel off the event horizon exponentially. For an arbitrary $\bar{v}_{0}$, we define $\hat{u}(\bar{u})$ by $\hat{u}(\bar{u}):=r(\bar{u},\bar{v}_{0})-r_{+}$, so that the solution of (130) can be written as
$r(\hat{u},\bar{v})\simeq r_{+}+\hat{u}e^{\kappa(\bar{v}-\bar{v}_{0})}.$ (131)
Note that the coordinate $\bar{u}$ in the metric (91) diverges at the horizon, but by construction $\hat{u}$ is regular.
A neighbourhood of the event horizon in asymptotically flat charged collapse at sufficiently late time (large $\bar{v}$) will be approximated by the event horizon in RN. As both $\hat{u}$ and our numerical coordinate $u$ are regular through the horizon, to leading order we must have $\hat{u}\simeq\hat{c}(u-u_{h})$ for some constant $\hat{c}>0$, where the horizon is at $u=u_{h}$. Differentiating (131) with respect to $u$, we find
$\Xi R\simeq cue^{\kappa\bar{v}}$ (132)
near the horizon at late times, where $c$ absorbs $\hat{c}$ and $\bar{v}_{0}$.
We have so far defined $\bar{v}$ only on RN, not on the collapse spacetime that we are trying to model. We note that on RN $\bar{v}$ obeys $\bar{v}=\bar{u}+2R$. If we assume that for sufficiently early $u$ and sufficiently late $\bar{v}$ the collapse spacetime is approximated by the RN exterior, we can characterize $\bar{v}$ on the collapse spacetime as the retarded time that obeys $\left.R_{,\bar{v}}\right|_{u}=2$ in that region. We could then define $\bar{v}=2R$ on $u=0$ and propagate it with $\Xi\bar{v}=0$.
Appendix E Boundary conditions on an ingoing null cone intersecting an outgoing one
We want to impose consistent null data on $x=0$, $u>0$ (the left boundary) and $u=0$, $x>0$ (the right boundary), intersecting at the point $u=x=0$ (the corner). On the left boundary we choose the gauge $B=0$ ($x=0$ is null) and $G_{,x}=0$ ($u$ is an affine parameter), and the electromagnetic gauge $A=0$. Then $\Xi=\partial_{u}$ and ${\cal H}=0$ there. In the following, we write $R(u,0)=:R_{l}(u)$ and $dR_{l}/du=:\dot{R}_{l}(u)$, and similarly for all other quantities on the left boundary.
The ingoing Raychaudhuri equation (48) with $A=B={\cal H}=0$ gives
$\ddot{R}_{l}+4\pi R_{l}(\dot{\psi}_{l}^{2}+\dot{\chi}_{l}^{2})=0.$ (133)
Note that $\ddot{R}_{l}\leq 0$, so if $\dot{R}_{l*}<0$ at the corner (as we shall always assume), then $\dot{R}_{l}<0$ for all $u\geq 0$. Eq. (28) with $A=B=0$ gives
$\dot{Q}_{l}={4\pi}qR_{l}^{2}(\chi_{l}\dot{\psi}_{l}-\psi_{l}\dot{\chi}_{l}),$ (134)
and Eq. (33) with $B=0$ gives
${2\dot{R}_{l}R_{,xl}\over G_{l}}=C_{l}-1,$ (135)
where $C_{l}$ is used as a shorthand for
$C_{l}:={2\over R_{l}}\left({\cal M}_{l}-{Q_{l}^{2}\over 2R_{l}}\right),$ (136)
compare the definitions (31) and (36). With $A=B=0$, and using (135) to eliminate $R_{,xl}$, (38) becomes
$\dot{\cal M}_{l}$ $=$ $-2\pi{R_{l}^{2}\over\dot{R}_{l}}(C_{l}-1)(\dot{\psi}_{l}^{2}+\dot{\chi}_{l}^{2})$ (137)
$+4\pi qQ_{l}R_{l}(\chi_{l}\dot{\psi}_{l}-\psi_{l}\dot{\chi}_{l}).$
Note that $\dot{\cal M}_{l}$ can have either sign. Complete boundary data are now obtained as follows.
[…]
${2\dot{R}_{l*}R_{,x*}\over G_{*}}=C_{*}-1.$ (138)
2) We fix the gauge on the right boundary $u=0$ by either setting $G(x,0)=G_{*}$ (affine initial gauge) or $R(x,0)=R_{*}+R_{,x*}\,x$ (Bondi initial gauge). We fix the gauge on the left boundary by setting $G_{l}(u)=G_{*}$ (affine left gauge). We now fix scalar field data $\psi_{l}(u)$, $\chi_{l}(u)$ on the left boundary and $\psi(0,x)$, $\chi(0,x)$ on the right boundary.
For the solution to be sufficiently often differentiable in a neighbourhood of the corner, we need to impose a sufficient number of corner conditions on the derivatives of the boundary data. Here we avoid this complication by assuming that the scalar field data are bounded away from the corner.
[…]
4) From (135) we read off
$V_{l}={C_{l}-1\over 2\dot{R}_{l}}.$ (139)
We now have the data that we need to start up the integration of the hierarchy equations at $x=0$, namely $Q=Q_{l}$ for (27), $A=0$ for (30), either $R=R_{l}$ and $V=V_{l}$ for (64,65) or $G=G_{*}$ for (39), $\Xi R=\dot{R}_{l}$ for (42), $\hat{\Xi}\psi=\dot{\psi}_{l}$ for (43), $\hat{\Xi}\chi=\dot{\chi}_{l}$ for (44), and ${\cal H}=0$ for (45).
In the simple case where $\psi_{l}=\chi_{l}=0$ the boundary ODE system becomes trivial, giving $Q_{l}=Q_{*}$, ${\cal M}_{l}={\cal M}_{*}$, and $R_{l}=R_{*}+\dot{R}_{l*}\,u$.
Appendix F Discretization of the left boundary equations
In the continuum, the data on $x=0$ are automatically compatible with the evolution equations for $\psi$, $\chi$ and $R$ and/or $G$ on $x\geq 0$. To make the discretised versions compatible as well, rather than $\psi_{l}$ and $\chi_{l}$ we specify $\dot{\psi}_{l}$ and $\dot{\chi}_{l}$, together with corner values $\psi_{*}$ and $\chi_{*}$. We write (40) in first-order form, and use the same ODE solver that we use to evolve $\psi$, $\chi$ and $R$ and/or $G$ in $u$ in the bulk to also evolve $Q_{l}$, ${\cal M}_{l}$, $\dot{R}_{l}$, $R_{l}$, $\psi_{l}$ and $\chi_{l}$ on the left boundary. With $A=B=0$ on the left boundary, the values of $\psi_{,u}$, $\chi_{,u}$, $R_{,u}$ and/or $(\ln G)_{,u}=0$ are then the same, up to round-off error, in the boundary ODE system and at $x=0$ in the bulk PDE system.
The Runge-Kutta ODE solver that we use for the time integration assumes that the ODE system is autonomous, we add $u$ as an evolved variable to the set of left-boundary variables, in order to access the given functions $\dot{\psi}_{l}(u)$ and $\dot{\chi}_{l}(u)$ at the correct substep values of $u$.
Appendix G Convergence testing
For any computed variable $f$, we define
${\cal E}[f](u,x;h):=\left({h\over h_{\text{ref}}}\right)^{-2}\left(f(u,x;h)-f(u,x;h/2)\right)$ (140)
where $h:=\Delta x=x_{\text{max}}/N_{x}$, and we assume $\Delta u$ is proportional to $h$. For a numerical quantity at resolution $h$, given that our numerical methods are designed to be second-order accurate, we expect the Richardson expansion
$f(u,x;h)=f_{0}(u,x)+h^{2}f_{2}(u,x)+h^{3}f_{3}(u,x)+O(h^{4}),$ (141)
where $f_{0}(u,x)$ is the (unknown) continuum solution and $f_{i}(u,x)$ are (discretisation-dependent, unknown) error terms, and hence
${\cal E}[f](u,x;h)=h_{\text{ref}}^{2}\left({3\over 4}f_{2}(u,x)+{7\over 8}hf_{3}(u,x)+O(h^{2})\right).$ (142)
Hence if ${\cal E}[f](u,x;h)$ for different $h$ are almost equal, we see pointwise second-order convergence, in the sense that the numerical error is dominated by the term $h^{2}f_{2}$. ${\cal E}[f](u,x;h)$ evaluated at any $h$ then provides an estimate of the actual error at the reference resolution $h=h_{\text{ref}}$.
We can also test the convergence (to zero) of $f:=M-\tilde{M}$, or of the difference $f$ of any variable computed in two different formulations. We then plot
${\cal E}[f](u,x;h):=\left({h\over h_{\text{ref}}}\right)^{-2}f(u,x;h),$ (143)
which, as $f_{0}(u,x)=0$, we expect to take values
${\cal E}[f](u,x;h)=h_{\text{ref}}^{2}\left(f_{2}(u,x)+hf_{3}(u,x)+O(h^{2})\right).$ (144)

### v2

@@
 I Introduction
-I.1 The problem posed by evolving through an event horizon in null coordinates
-In this series of papers, starting with axinull_formulation , we have been developing numerical simulations of gravitational collapse on null cones emanating from a regular centre: in the present paper, we examine formulations that allow us to evolve from regular data through the black hole event horizon.
-After this Introduction, we will restrict to spherical symmetry in 3+1 spacetime dimensions, but the ideas generalise straightforwardly, and will be applied to non-spherical critical collapse elsewhere.
+I.1 Overview
+In this series of papers, starting with axinull_formulation , we have been developing numerical simulations of gravitational collapse on null cones emanating from a regular centre.
+Null coordinates are an established tool in spherical symmetry, but our main motivation is simulating critical collapse on null cones beyond spherical symmetry.
+As long as no caustics form, null coordinates are ideal for type-II critical collapse because they allow us to adapt the coordinate system to the approximate self-similarity and large range of scales seen there.
+In the present paper, we introduce new formulations of the Einstein equations in null coordinates that allow us to evolve from regular data through the black hole event horizon, while keeping the radial coordinate arbitrary.
+The motivation for this will be given in more detail in the next Subsection, but, in brief, we want to evolve through horizons to accurately characterise black holes, while we also want to keep the radial coordinate arbitrary in order to adapt the coordinates to type-II critical collapse.
+To make the presentation more concrete, after this Introduction we apply the new formulations to the spherical Einstein-Maxwell-charged scalar field system.
+After demonstrating convergence, we turn to type-II critical collapse in this system, and obtain some new physis results.
+In particular, we verify the critical exponents for the black hole mass and charge, calculate the mass and charge fine structures, and present preliminary evidence that the charge fine structure is not universal.
+We plan to return to non-spherical type-II critical collapse elsewhere.
+The spherical Einstein-Maxwell-charged scalar systems is of current interest also because of the stable extremal critical collapse conjecture of Kehle and Unger KehleUnger22 ; KehleUnger24 .
+This says that there is an open region of solution space in which the threshold of black hole formation consists of extremal black holes.
+In this context, we will need to evolve initial data posed on an outgoing and an ingoing null cone that meet at 2-sphere.
+The domain of dependence of these data is then a “null rectangle” in the symmetry-reduced spacetime.
+We present an algorithm for this setup, and demonstrate convergence in selected example solutions.
+We plan to investigate extremal critical collapse in this system elsewhere.
+I.2 The problem posed by evolving through an event horizon in null coordinates
 In Paper II of the series axinull_critscalar , we investigated the threshold of black hole formation for a massless scalar field coupled to general relativity in axisymmetry, adapting the coordinates to the expected self-similar critical solution.
-While, at moderate non-sphericity, this allowed us to fine-tune to the threshold of black hole formation to machine precision, without any need for adaptive mesh-refinement, we could not reach non-sphericities as high as others had reached before us.
+While, at moderate non-sphericity, this allowed us to fine-tune to the threshold of black hole formation down to machine precision, without any need for adaptive mesh-refinement, we could not reach non-sphericities as high as others had reached before us.
 In hindsight, this was due both to our choice of radial gauge (how the radial coordinate $x$ labels points on each null cone generator) and our choice of formulation (which of the Einstein equations are used in the code and which are not).
@@
 Here we evolve $G$, which allows us to use an arbitrary radial shift $B$, and in particular one adapted to critical collapse.
-We also present a third formulation where both $R$ and $G$ are evolved, and the Raychaudhuri equation is not used.
-In the literature, this has been used only with double null gauge $B=0$, but $B$ can in fact be chosen freely.
-Here we point out that all three formulations can be combined with any radial gauge choice.
-We present extended tests of these three formulations in spherical symmetry, with charged scalar field matter.
+We also present a third formulation, where both $R$ and $G$ are evolved and the Raychaudhuri equation is used only to construct initial data.
+In the literature, this formulation has been used only with double null gauge $B=0$, but $B$ can in fact be chosen freely.
+One of our main points is the observation that all three formulations can be combined with any radial gauge choice.
 The gauge problem in critical collapse mentioned above does not arise in spherical symmetry, and will be addressed when we return to axisymmetry elsewhere.
-I.2 Formulations and radial gauges in null coordinates
-We now give a more precise statement of what we just said about formulations, but without restricting the spacetime dimension or symmetry.
-(In the remainder of the paper, we will restrict to spherical symmetry in 3+1 dimensions.) We remind the reader that we consider line elements of the form
+I.3 Formulations and radial gauges in null coordinates
+We now give a more precise statement of what we just said about formulations.
+For clarity, we do not fix the spacetime dimension or symmetry.
+(In the remainder of the paper, we will restrict to spherical symmetry in 3+1 dimensions.) We consider line elements of the form
 $ds^{2}$ $=$ $-2G\,du\,(dx+B\,du)$ (1)
 $+R^{2}\gamma_{ij}(d\theta^{i}+\beta^{i}\,du)(d\theta^{j}+\beta^{j}\,du).$
-(The quantity $B$ replaces the metric coefficient $H:=-g_{uu}$ that we used in Papers I–III.
-This is just for convenience of notation.) The spacetime dimension is $n+2$, with the $\theta^{i}$ coordinates on an $n$-dimensional manifold, typically $S^{n}$.
+(The quantity $B$ replaces the metric coefficient $H:=-g_{uu}$ that we used in Papers I–III, with $B:=H/(2G)$.
+This is just for convenience of notation.) The spacetime dimension is $n+2$, and the $\theta^{i}$ are coordinates on an $n$-dimensional manifold, typically $S^{n}$.
 From this form of $g_{\mu\nu}$ it follows that $g^{uu}=g^{ui}=0$, and $g^{ux}=-G^{-1}$.
@@
 We must have $G>0$ for the metric to be regular.
-$R$ is defined by restricting $\gamma_{ij}$ to have a fixed determinant (such as its value in Minkowski spacetime).
+$R$ is defined by restricting $\gamma_{ij}$ to have a fixed determinant (such as its value on a unit sphere $S^{n}$ in standard coordinate $\theta^{i}$).
 The remaining gauge freedom in this ansatz for the line element is to re-label $u$ as $u\to u^{\prime}(u)$, and to relabel the radial coordinate $x$ as $x\to x^{\prime}(u,x,\theta^{i})$.
@@
 ${d\over d\lambda}:=G^{-1}\partial_{x}.$ (4)
-(See also Paper I axinull_formulation for the full definition of $\rho_{+}$.) The quantity
+(See also Paper I axinull_formulation for the definition of $\rho_{+}$ from first principles.) The quantity
 $\lambda(u,x,\theta^{i}):=\int_{0}^{x}G(u,x^{\prime},\theta^{i})\,dx^{\prime}$ (5)
@@
 where the term $K$ contains shear and matter terms, and is non-negative in vacuum or with matter obeying the dominant energy condition.
-From (8), with $K/G\geq 0$, we see that once $dR/d\lambda$ has become negative, $R$ reaches zero in finite affine parameter.
+From (8), with $K\geq 0$, we see that once $dR/d\lambda$ has become negative, $R$ reaches zero in finite affine parameter.
 Equivalently, from (9) we see that once $\rho_{+}$ has become negative, it blows up in finite affine parameter.
 We can now make the above statement precise: In a non-spherical collapse spacetime, generically there will be $u_{0}<u_{1}$ such that for any coordinate null cone with $u_{0}<u<u_{1}$, the expansion $\rho_{+}$ remains positive on some of its generators, while on others it becomes negative and then blows up.
-This could be because it has reached a curvature singularity, or because the null cone has developed a caustic.
+This could be because the generator has reached a curvature singularity, or because the null cone has developed a caustic.
 We now come to the role of the Raychaudhuri equation in numerical relativity on null cones.
-In its form (6) and with $G$ given, it can be solved as a second-order ordinary differential equation (ODE) in $x$ for $R$.
+In its form (6) and with $G$ given, it can be solved as a second-order ordinary differential equation (from now on, ODE) in $x$ for $R$.
 In its form (7) and with $R$ given, it can be solved as a first-order ODE for $G$.
@@
 The observation that the choices of formulation and gauge are independent is one of the main points of the present paper.
+One can retain some of the stability advantages of a maximally constrained formulation in the eR formulation by restricting to the family of “local shifted Bondi” (lsB) radial gauges, where $R=R(u,x)$.
+We did this in Paper II.
+Similarly, in the eG formulation, we can restrict to “local shifted affine” (lsa) gauges, where $G=G(u,x)$, to make it almost maximally constrained.
 Formulation Initial gauge Evolved gauge Restricted to Gauge evolution equations
@@
 Table 1: Summary of the five formulations discussed in Sec.
-I.2.
+I.3.
 “Initial gauge” denotes the freedom to label points on $u=0$ by the coordinate $x$, so “any RR” is short for “any choice of $R(0,x,\theta^{i})$” and similarly for “any GG”.
@@
 In our own previous work, we have explored the following additional combinations:
-I.3 Plan of this paper
-In the following, we will restrict to spherical symmetry in 3+1 dimensions, with charged scalar field matter, but the ideas illustrated here are easily generalised to less symmetry, other spacetime dimensions, and to vacuum or other matter.
-We present an implementation of all three formulations set out above in one code, and test them in two settings.
+I.4 Plan of this paper
+In the remainder of this paper, we will restrict to spherical symmetry in 3+1 dimensions, with charged scalar field matter, but the ideas illustrated here are in principle easily generalised to less symmetry, other spacetime dimensions, and to vacuum or other matter.
+We present an implementation of the five formulations described above in one code, and we test the three formulations that admit a general radial gauge in two settings.
 In the first of these, we evolve on outgoing null cones that emanate from a regular central worldline.
 To fix the radial gauge in the initial data, we either set $R=x$ or $G=1$ at $u=0$.
-We continue in “shifted double-null gauge”, an incremental gauge that is adapted to critical collapse in spherical symmetry (see Sec.
+We continue in “shifted double-null gauge”, see Eq.
+(60) below, an incremental gauge that is adapted to critical collapse in spherical symmetry (see also Sec.
 III.C of Paper I for a definition and PortoGundlach2022 for a previous application).
@@
 We continue the initial gauge choice of $R_{x,}=1$ or $G=1$ as double-null gauge $B=0$.
+We do not present any results in the affine and Bondi formulations, as the affine and Bondi gauges that they require do not allow us to make the outer boundary future spacelike (in the first setting) or ingoing null (in the second setting).
 In both settings, we present evidence of second-order convergence with resolution, both within each formulation and between the three pairs of formulations.
@@
 VII.
-II Field equations
+II The Einstein-Maxwell-charged scalar system
 II.1 Field equations in covariant form
@@
 $+{{1\over 4\pi}}\left(F_{ac}{F_{b}}^{c}-{1\over 4}g_{ab}F_{cd}F^{cd}\right).$
-The equations admit the gauge freedom $\phi\to e^{iq\alpha}\phi$, $A_{a}\to\nabla_{a}\alpha$ for an arbitrary scalar function $\alpha$, leaving $F_{ab}$ and $D_{a}\phi$, and hence $j^{a}$ and $T_{ab}$, invariant.
+The equations admit the gauge freedom $\phi\to e^{iq\alpha}\phi$, $A_{a}\to A_{a}+\nabla_{a}\alpha$ for an arbitrary scalar function $\alpha$, leaving $F_{ab}$ and $D_{a}\phi$, and hence $j^{a}$ and $T_{ab}$, invariant.
 II.2 Field equations in spherical symmetry in null coordinates
 We consider the general metric in spherical symmetry with one null coordinate $u$,
-$ds^{2}=-2G\,du(dx+B\,du)+R^{2}\,d\Omega^{2}.$ (19)
+$ds^{2}=-2G\,du(dx+B\,du)+R^{2}\,d\Omega^{2},$ (19)
 where $G$, $B$ and $R$ are functions of $(u,x)$, and $d\Omega^{2}:=d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}$.
@@
 can be achieved by the gauge transformation with
-$\alpha(u,x)=\int_{0}^{u}A_{u}(u^{\prime},0)\,du^{\prime}-\int_{0}^{x}A_{x}(u,x^{\prime})\,dx^{\prime}.$ (21)
+$\alpha(u,x)=-\int_{0}^{u}A_{u}(u^{\prime},0)\,du^{\prime}-\int_{0}^{x}A_{x}(u,x^{\prime})\,dx^{\prime}.$ (21)
 From now on, we work exclusively in this gauge, and for conciseness we rename $A_{u}$ to $A$. Note that $A_{x}=0$ can be written geometrically as $U^{a}A_{a}=0$.
@@
 $C(u,x)=1+{2R_{,x}\Xi R\over G}.$ (33)
-As $R\geq 0$ and $G>0$, $\rho_{+}$ has the same sign as $R_{,x}$.
-Hence a symmetry sphere is trapped if and only if $R_{,x}<0$, and marginally trapped if and only if $R_{,x}=0$.
-(By trapped, we always mean outer-trapped.)
 The derivatives of the Misner-Sharp mass are
@@
 $+{4\pi}qQR(\chi\hat{\Xi}\psi-\psi\hat{\Xi}\chi).$
-Hence there is another conserved mass curent analogous to (29).
+Hence there is another conserved mass current analogous to (29).
 […]
@@
 which can be integrated for $G$, given $R$, $\psi$ and $\chi$, but only if $R_{,x}>0$. Alternatively, we can write it as
-$R_{,xx}-{G_{,x}\over G}R_{,x}+4\pi R(\psi_{,x}^{2}+\chi_{,x}^{2})=0,$ (40)
+$R_{,xx}-{G_{,x}\over G}R_{,x}+4\pi(\psi_{,x}^{2}+\chi_{,x}^{2})R=0,$ (40)
 which can be solved as a second-order linear ODE for $R$, given $G$, $\psi$ and $\chi$. There is then no restriction on the sign of $R_{,x}$. More geometrically, $E_{xx}=0$ can be written as
@@
 In double-null gauge $B=0$, we have $\bar{G}=G$ and $X=G^{-1}\partial_{u}$, in analogy to $U=G^{-1}\partial_{x}$ (the latter holds in any radial gauge).
-In initial data on an outgoing null cone with regular centre, we always have $\Xi R<0$ and hence $XR<0$ on our initial data.
+In initial data on an outgoing null cone with regular centre, we always choose $\Xi R<0$ and hence $XR<0$ on $u=0$.
 From (49) we see that $XXR\leq 0$, so $XR<0$ and hence $\Xi R<0$ on the entire domain of dependendence of these initial data.
-Hence, from (33), a symmetry sphere is trapped if and only if $C>1$ and marginally trapped if and only if $C=1$.
-This is not obvious from (42), whose right-hand side becomes negative when $R<|Q|$, and so, as we solve (42) but not (49) in algorithm, $\Xi R<0$ may be violated by numerical error.
+This is not obvious from (42), whose right-hand side becomes negative when $R<|Q|$, and so, as we solve (42) but not (49), $\Xi R<0$ may be violated by numerical error.
+As $R\geq 0$, $G>0$ and $\Xi R<0$, $\rho_{+}$ given above in (3) has the same sign as $R_{,x}$. Hence a symmetry sphere is outer-trapped if and only if $R_{,x}<0$, or equivalently $C>1$.
 The choice of coordinate $x$ for $u>0$ is controlled incrementally by the choice of $B$.
@@
 To see this, note that if we make a transformation from double-null coordinates $(u,v)$ to general null coordinates $(u,x)$, parameterised by $v=v(u,x)$, then
-$-2\bar{G}\,du\,dv$ $=$ $-2G\,du(dx+B\,du),$ (53)
+$2\bar{G}\,du\,dv=-2G\,du(dx+B\,du)$ (53)
+holds if and only if
 $G$ $=$ $\bar{G}v_{,x},$ (54)
 $v_{,u}$ $=$ $Bv_{,x},$ (55)
-regardless of spacetime dimension or symmetry.
+regardless of spacetime dimension or symmetry, and so $v$ is advected in $x$ with speed B-B.
 […]
@@
 III Algorithm
+s
 III.1 Initial data and gauge choice
@@
 III.2 Boundary conditions at $x=0$
-Where $R=x=0$ is a regular central worldline, $x=0$ is a coordinate singularity (and we omit it from the numerical grid), so we cannot impose simple boundary conditions there.
+When $x=0$ is the regular central worldline $R=0$, $x=0$ is a coordinate singularity, and we cannot impose simple boundary conditions there.
 To start up the integration of the hiercharchy equations, we expand in powers of $x$ instead.
 At a given $u$, we fit the evolved quantities $\psi$, $\chi$, $R/x$ and/or $G$ either to linear functions of $x$, such as $\psi\simeq\psi_{(0)}+\psi_{(1)}x$, or to quadratic ones, such as $\psi\simeq\psi_{(0)}+\psi_{(1)}x+\psi_{(2)}x^{2}$.
-(The notation here is compatible with our notation $\psi_{l(n)}(u)$ in Papers I and II, where $l$ denotes the spherical harmonic and $n$ the power of $x$.)
+(The notation here is compatible with our notation $\psi_{l(n)}$ in Papers I and II, where $l$ denotes the spherical harmonic and $n$ the power of $x$.)
 $R_{(1)}$, $G_{(0)}$ and ${\cal H}_{(0)}$ are not determined by the equations.
 However, we show in Appendix B that at a regular centre, and with $u$ proper time at the centre, they obey $G_{(0)}=R_{(1)}$, $B_{(0)}=1/(2R_{(1)})$ and ${\cal H}_{(0)}=0$.
-The full expressions are given in Appendix C.
-On the other hand, if $x=0$ is a timelike or null inner boundary, we can specify boundary conditions for the hierarchy equations there, subject to constraints.
+The full expressions for the expansion coefficients are given in Appendix C.
+On the other hand, if $x=0$ is a timelike or null boundary, we can specify boundary conditions for the hierarchy equations there, subject to constraints.
 For the case where $x=0$ is an ingoing null cone, we discuss this in detail in Appendix E.
 III.3 Time evolution algorithm
-We use the “method of lines” framework, where one discretises separately in space and time. We break up each sub-timestep into the following steps, where Steps 1 and 3 depends on the formulation.
-Step 0: Assume that on a null cone of constant $u$, the functions $\psi(x)$, $\chi(x)$, and $\ln G(x)$ and/or $R(x)$ are given.
+We use the “method of lines” framework, where one discretises separately in space and time.
+We break up each sub-timestep into the following steps, where Steps 1 and 3 depend on the formulation.
+For completeness, we include the affine and Bondi formulations in the following pseudo-code, even though we do not use them for any numerical tests in this paper.
+Step 0: Assume that on a null cone of constant $u$, the functions $\psi(x)$, $\chi(x)$, and $G(x)$ and/or $R(x)$ are given.
 When $G$ and $R$ are both given, they must obey the constraint (40).
@@
 Step 1b: Alternatively, in the eG or affine formulation, we solve the Raychaudhuri equation in its form (40) as a homogeneous linear second-order ODE for $R$, and we set the fitting coefficients of $R$ consistently.
-Step 1c: Alternatively, in free evolution, we do nothing.
+Step 1c: Alternatively, in the fe formulation we do nothing.
 From numerical error, both in the evolution of $R$ and $G$, and in the fitting, the fitting coefficients of $R$ and $G$ will not be exactly compatible with the Raychaudhuri equation, so we need to make an arbitrary choice which set of fitting coefficients we choose for expanding the solution of the hierarchy equations at the centre.
@@
 Step 3b: Alternatively, in the Bondi formulation, $B$ is given explicitly by (57).
-Step 3c: Alternatively, in the other formulations we set $B$ freely. For critical collapse in spherical symmetry, we use sdn gauge (60). On a null rectangle we use double-null gauge $B=0$.
+Step 3c: Alternatively, in the eR, eG and fe formulations we set $B$ freely. For critical collapse in spherical symmetry, we use sdn gauge (60). On a null rectangle we use double-null gauge $B=0$.
 Step 4: We find the $u$-derivatives of $\psi$, $\chi$, $R$ and/or $\ln G$ from their $\Xi$-derivatives and $x$-derivatives, using (22).
@@
 At $x=0$ this makes the upwind derivative a right derivative, and at $x_{\text{max}}>x_{0}$ a left derivative.
-Step 5: Using the method of lines, we evolve $\psi$, $\chi$, and one or both of $R$ and $\ln G$, from $u$ to $u+\Delta u$. This returns us to Step 0.
+Step 5: Using an ODE integrator at each grid point (the method of lines), we evolve $\psi$, $\chi$, and one or both of $R$ and $\ln G$, from $u$ to $u+\Delta u$. This returns us to Step 0.
 IV Numerical methods
@@
 $(f_{,x})_{i-{1\over 2}}$ $\simeq$ ${f_{i}-f_{i-1}\over\Delta x}.$ (63)
-This scheme is by construction second-order accurate, and is causal and explicit in the sense that the vector of variables $g_{i}$ depends only on $g_{j}$ with $j<i$ and $f_{j}$ with $j\leq i$: information propagates along a null cone of constant $u$ only towards increasing $x$.
+This scheme is explicit, is second-order accurate by construction, and is causal in the sense that the vector of variables $g_{i}$ depends only on $g_{j}$ with $j<i$ and $f_{j}$ with $j\leq i$: information propagates along a null cone of constant $u$ only towards increasing $x$.
 To solve the Raychaudhuri equation (40) for $R$, we write it in the first-order form
+$R_{,x}$ $=$ $G\,V,$ (64)
 $V_{,x}$ $=$ $-{K\over G}\,R,$ (65)
-$R_{,x}$ $=$ $G\,V,$ (65)
 with
@@
 $K$ $:=$ $4\pi(\psi_{,x}^{2}+\chi_{,x}^{2}).$ (67)
-When we solve (64,65) for $(V,R)$, the coefficients $G$ and $K$ are already known.
+When we solve (65,64) for $(V,R)$, the coefficients $G$ and $K$ are already known.
 $K$ is naturally given at midpoints, but $G$ needs to be interpolated there using (62). We discretize this system, using a 2-point stencil, to second-order accuracy as
@@
 Hence we can expect the numerical error to be reasonably small.
-For comparison, we also use the second or fourth-order Adams-Bashforth method. These need the vector of right-hand sides $F_{i}$ at grid points. We discrete derivatives on grid points as
+For comparison, we also use the second or fourth-order Adams-Bashforth method. These need the vector of right-hand sides $F_{i}$ at grid points. We discretize $x$-derivatives on grid points as
 $(f_{,x})_{i}\simeq{f_{i+1}-f_{i-1}\over 2\Delta x}$ (73)
 where this is defined, and the standard one-sided three-point stencils at the left and right boundaries $i=1$ and $i=N_{x}$. This is also second-order accurate but not causal.
-When we solve (39) for $G$, we first find $\ln V$, where $V:=R_{,x}/G$, on grid points by integration.
+When we solve (39) for $G$, we first find $\ln V$ on grid points by integration, where $V:=R_{,x}/G$.
 (In Papers I-III, we used the variables $g:=1/V$ and $\gamma:=\ln g$.
@@
 To evaluate the compactness $C$ when $G$ is constrained, we use $V$ on grid points.
-As $G$ varies rapidly in $u$ and $x$, but is constrained to be strictly positive, we evolve $\ln G$ rather than $G$.
+$G$ varies rapidly in $u$ and $x$, but is constrained to be strictly positive.
+To avoid it becoming negative through numerical error, we evolve $\ln G$ rather than $G$.
 We find $G$ at grid points by exponentiating, and hence on mid-points by linear interpolation.
@@
 From its definition, the quantity $Q/R^{3}$, representing an average charge density, is expected to be smooth, and to be non-zero at the centre.
-Numerically, this is of course not guaranteed.
+Numerically, this smoothness is not guaranteed.
 We can improve regularity by discretizing
@@
 $\Delta u_{2}:=C_{2}\left(\max_{x}\left|B\right|\right)^{-1}\,\Delta x$ (78)
-for some $C_{2}\lesssim 1$. In all our gauges $R_{,u}=0$ at $R=0$, and so (77) and (78) are equivalent at the centre.
+for some $C_{2}\lesssim 1$. In all our gauges $R_{,u}=0$ at $R=0$, and so (77) and (78) with $C_{1}=C_{2}$ are equivalent at the centre.
 Since Paper II, we have realised that using $\Delta u=\Delta u_{1}$ may be an unnecessarily small timestep.
-We have found that $\Delta u=\Delta u_{2}$ with $C_{1}=0.1$ also leads to stable evolutions up to close to the apparent horizon, but with fewer timesteps overall.
+We have found that $\Delta u=\Delta u_{2}$ with $C_{2}=0.1$ also leads to stable evolutions up to close to the apparent horizon, but with fewer timesteps overall.
 As the apparent horizon is approached, $\Delta u_{1}\to 0$, and so this timestep does not allow us to evolve through the horizon.
@@
 The time steps $\Delta u_{3}$ and $\Delta u_{4}$ cannot be used for convergence testing as $\Delta u_{3}$ is independent of $\Delta x$.
-As an alternative to all these, for example if we want to test convergence of a solution where $R_{,x}=0$ occurs, we can also use the simple timestep
+As an alternative to all these, for example if we want to test convergence of a solution in double-null gauge where $R_{,x}=0$ occurs, we also use the simple fixed timestep
 $\Delta u_{0}:=C_{0}\Delta x.$ (81)
 IV.4 Singularity excision
-As no influence propagates from smaller to larger $x$ in our code for all $x$ with $B<0$, and we are in spherical symmetry, we can use a crude type of similarity excision where we excise gridpoints $i\geq i_{\text{exc}}$, where $i_{\text{exc}}$ is the first grid point where either $R_{i}<0$ or $R_{i}<R_{i-1}/2$.
+In the integration of the hierarchy equations from $x=0$, information propagates only from smaller to larger $x$.
+The same is true for the time evolution equations (74), whereever $B<0$.
+Our discretisation of the hierarchy and evolution equations respects this causality.
+In spherical symmetry, we can therefore use a crude type of singularity excision where we excise gridpoints $i\geq i_{\text{exc}}$, where $i_{\text{exc}}$ is the first grid point where either $R_{i}<0$ or $R_{i}<R_{i-1}/2$.
 (The second criterion makes sure that the first non-excised point is not too close to $R=0$.) It is essential that we re-determine $i_{\text{exc}}$ after each time we have solved (40) for $R$, that is, in all sub-timesteps.
-Obviously $i_{\text{exc}}$ is not allowed to decrease from one sub-timestep to the next, or one timestep to the next.
+Obviously $i_{\text{exc}}$ is not allowed to decrease from one sub-timestep to the next, or from one full timestep to the next.
 Once excision starts, the excision radius $x_{\text{exc}}(u):=i_{\text{exc}}(u)\Delta x$ decreases quickly to zero.
@@
 With a charged scalar field, $q$ sets an inverse length scale in the field equations, so $qx_{c}$, $qd$ and $q/\omega$ are all relevant dimensionless parameters.
-This also means that changing $q\to\lambda q$ is equivalent to $x_{c}\to x_{c}/\lambda$, $d\to d/\lambda$, and similarly for all other length scales in the initial data and $\omega\to\lambda\omega$.
+This also means that changing $q\to\lambda q$ is equivalent to $x_{c}\to x_{c}/\lambda$, $d\to d/\lambda$, and $\omega\to\lambda\omega$, and similarly for any other length scales in the initial data.
+Therefore, it may be convenient to fix a length scale in the initial data, such as one of $x_{c}$, $d$ and $\omega$ in (82), and consider $q$ as a parameter of the initial data in its place (rather than a constant of nature).
+Thus, in one test below, we keep $x_{c}$, $d$ and $\omega$ the same and vary $q$.
+Varying $q$ has the conceptual advantage that for given initial data $\psi(x)$, $\chi(x)$ and $R(x)$ or $G(x)$, $Q(x)$ in the initial data is manifestly proportional to $q$.
 V.2 Convergence tests
-For convergence testing, we have focused on initial data (82) with $x_{c}=0.5$, $d=0.5$, $\omega=2.0$, and the coupling constant $q=1/\sqrt{4\pi}$.
+For convergence testing, we have focused on the initial data (82) with $x_{c}=0.5$, $d=0.5$, $\omega=2.0$, and the coupling constant $q=1/\sqrt{4\pi}$.
 We complete the initial data either with $G=1$, and then set $p{\cal A}=0.15$, or with $R=x$, and then set $p{\cal A}=0.2$.
@@
 This is not in itself a problem, as $h^{2}f_{2}(x_{i})\simeq h^{2}/(ih)=h/i$, so even at $i=1$ the error goes down as $h$.
-However, we find that by expanding $R\hat{\Xi}\psi$ and $R\hat{\Xi}\psi$ to second order, we can essentially remove the $h^{2}f_{2}$ error, so that the error is dominated by $h^{3}f_{3}$.
+However, we find that by expanding $R\hat{\Xi}\psi$ and $R\hat{\Xi}\chi$ to second order, we can essentially remove the $h^{2}f_{2}$ error, so that the error is dominated by $h^{3}f_{3}$.
 This reduces the actual error at $i=1$ by two or three orders of magnitude already at $h=0.04$, and by more at higher resolution.
@@
 […]
-We note in passing that at $u=0$, where $\psi$, $\chi$ and either $R$ or $G$ or both are still exact, $\Xi\psi$ and $\Xi\psi$ converge to third order.
+We note in passing that at $u=0$, where $\psi$, $\chi$ and either $R$ or $G$ or both are still exact, $\Xi\psi$ and $\Xi\chi$ converge to third order.
 The results for the six combinations of initial Bondi gauge or initial affine gauge with the eR, eG and fe formulations are broadly similar.
@@
 $E_{xx(1)}$ $:=$ $GV_{,x},$ (83)
-$E_{xx(2)}$ $:=$ $-4\pi R(\psi_{,x}^{2}+\chi_{,x}^{2})$ (84)
-using a different methods from elsewhere in the code.
+$E_{xx(2)}$ $:=$ $-4\pi R(\psi_{,x}^{2}+\chi_{,x}^{2}),$ (84)
+using a different method from elsewhere in the code.
 This means in the eR or eG formulations, where $G$ or $R$, respectively, are found by solving $E_{xx}=0$, we have a non-trivial test of the discretization of that process.
@@
 V.4 Type-II critical collapse of a spherical charged scalar field
-As our emphasis in this paper is on developing methods for simulating type-II critical collapse on null cones beyond spherical symmetry, rather than on examining charged collapse in spherical symmetry, we examined only two one-parameter families of initial data, namely (82) with $x_{c}=1.5$, $d=0.5$, $\omega=2.0$ and ${\cal A}=0.05$, for either $R(0,x)=x$ or $G(0,x)=1$.
-We set the coupling constant to $q=1/\sqrt{4\pi}$.
+As our emphasis in this paper is on developing methods for simulating type-II critical collapse on null cones beyond spherical symmetry, rather than on charged collapse in spherical symmetry in detail, we examine only four one-parameter families of initial data.
+In the first of these, $\phi(0,x)$ is given by (82) with $x_{c}=1.5$, $d=0.5$, $\omega=2.0$ and ${\cal A}=0.05$, and we initialise the radial gauge with $G(0,x)=1$.
+We set the coupling constant to $q=1/\sqrt{4\pi}\simeq 0.2821$.
+The second family is as the first except that we initialise the radial gauge with $R(0,x)=x$.
+As already $R_{,x}(0,x)\simeq 1$ in the first family, these two families do not differ qualitatively.
+The third family is again like the first, except that we set $\omega=0$, so that the scalar field is real and $Q=0$.
+The fourth family is again like the first, except that we set $q=10/\sqrt{4\pi}\simeq 2.821$, so that $Q(0,x)$ is 10 times as large.
 During the evolution, we use sdn gauge, given in (60), with the parameter $x_{0}$ adjusted so that $x=x_{0}$ is approximately the past light cone of the accumulation point of echos in near-critical solutions.
@@
 In the following, for conciseness we present results only for the $G=1$ family of initial data, evolved with the eG formulation.
-As a proxy for the black hole (Hawking) mass and charge, which are strictly speaking defined at future null infinity, we evaluate $M$ and $Q$ on the first symmetry 2-sphere where the Hawking compactness takes the value $C(u,x)=C_{\text{thr}}$.
+As a proxy for the black hole Hawking mass and charge, which are strictly speaking defined at future null infinity, we evaluate $M$ and $Q$ on the first symmetry 2-sphere where the Hawking compactness takes the value $C(u,x)=C_{\text{thr}}$.
 (“First” here means the smallest such $u$.) In the results presented here in the eG formulation, we choose the obvious value $C_{\text{thr}}=1$, but this is not possible in the eR formulation.
@@
 Figure 1: Maximum curvature fine-structure for the $G=1$ family of initial data, in the eG formulation: we plot $1/2\ln|\max T|-\gamma\ln(1-p/p_{*})+A$ against $\ln(1-p/p_{*})$.
-We use the approximate theoretical value $\gamma=0.374$, and the family-dependent offset A=0.23A=-0.23 has been fitted by eye so that the average of the fine-structure is approximately zero at high fine-tuning.
+We use the approximate theoretical value $\gamma=0.374$, and the family-dependent offset A=0.23A=-0.23 has been fitted by eye so that the average of the fine-structure is approximately zero at high fine-tuning to the threshold of collapse,(that is, in the limit $|p-p_{*}|\to 0$).
 The range of $\ln(1-p/p_{*})$ on the horizontal axes has been chosen as in Fig.
 1 of Paper II for comparison.
-(As in Paper II, we have suppressed axis labels to make figures larger.) The thin dotted line has been fitted by eye to the first two peaks from the right and has slope $\delta\gamma=0.012$, meaning that the curvature critical exponent is slightly larger at low fine-tuning.
+(As in Paper II, we have suppressed axis labels to make figures larger.) The thin dotted line has been fitted by eye to the first two peaks from the right and has slope $\delta\gamma=0.012$, meaning that the curvature critical exponent is slightly larger at low fine-tuning
 Figure 2: Black hole mass fine-structure for the $G=1$ family of initial data, in the eG formulation, with $C_{\text{thr}}=1$: we plot $\ln M-\gamma\ln(p/p_{*}-1)+A$ against $\ln(p/p_{*}-1)$.
@@
 The thin dotted line has been fitted by eye to the first two peaks from the right.
-It has slope $\delta\gamma=0.010$, meaning that the mass critical exponent is slightly larger at low fine-tuning, in line with that for that of the curvature to within a plausible fitting error.
+It has slope $\delta\gamma=0.010$, meaning that the mass critical exponent is slightly larger at low fine-tuning, in line with that for the curvature, to within a plausible fitting error.
 Figure 3: Black hole charge fine-structure for the $G=1$ family of initial data, in the eG formulation, with $C_{\text{thr}}=1$: we plot $\ln Q-\mu\ln(p/p_{*}-1)+A$ against $\ln(p/p_{*}-1)$.
@@
 On the collapse side, Figs.
-2 and 3 show the fine-structure for $M$ and $Q$ of the first (approximate) MOTS.
+2 and 3 show the fine-structure for $M$ and $Q$ of the first (approximate) marginally outer-trapped surface (from now on, MOTS and FMOTS for the first MOTS).
 What we actually plot is $1/2\ln T-\gamma\ln(1-p/p_{*})+A$ in Fig.
@@
 At high fine-tuning, the fine-structures are periodic, thus confirming the values of $\gamma$ and $\mu$ from perturbation theory.
-The fine structure for $T$ has the same amplitude of about $0.15$ and flat top as in Fig.
+The fine structure for $T$ has the same amplitude of about $0.15$ and shape as in Fig.
 1 of Paper II, which showed $T$ for a one-parameter family of real, spherically symmetric initial data.
 At low fine-tuning, the curve has the same periodic structure but rises slightly with $\ln(1-p/p_{*})$.
-The same observation applies to the periodic fine structure of $M$ at high fine-tuning.
+The same observation applies to the periodic fine structure of $M$.
 It has an amplitude $\simeq 0.03$.
@@
 3, the fine structure in $\ln Q$ evaluated also appears to be continuous, with amplitude $\simeq 0.15$.
-The value of $\mu$ predicted in GundlachMartin96 was verified in collapse simulations in HodPiran97 and Petryk05 , and is verified here again by us.
+The value of $\mu$ predicted in GundlachMartin96 was verified in collapse simulations in HodPiran97 and Petryk05 , and is verified here again by us to be correct sufficiently close to the threshold of collapse.
 Figures in these last two papers showed a power law with a wiggle, but did not show the wiggle clearly.
@@
 At low fine-tuning, we see deviations from the expected scaling laws.
-Heuristically, these can be described as modifications of the critical exponents $\gamma$ and $\mu$ and the echoing period $\Delta$ (as we did for the critical exponents for the mass and angular momentum in GunBau18 ).
+Heuristically, these can be described as modifications of the critical exponents $\gamma$ and $\mu$ (as we did for the critical exponents for the mass and angular momentum in GunBau18 ).
 To quantify these modifications, we have fitted the thin dotted lines in Figs.
@@
 As in GunBau18 , the small change in critical exponent is a convenient quantitative way of describing the fine structure at low fine-tuning, but we do not have a quantitative explanation for it.
-For comparison, we have also fine-tuned a family of uncharged $G=1$ initial data, differing from our charged data only by the choice $\omega=0$.
+For comparison, we have also fine-tuned a family of real, and therefore uncharged, initial data with $G=1$, differing from our charged data only by the choice $\omega=0$.
 We also set ${\cal A}=0.075$ in order to make $p_{*}\simeq 1$.
@@
 1, 2 and 3, this seems to be accurate for $\ln|1-p/p_{*}|\lesssim-20$.
-This corresponds to $|Q|/M\lesssim 10^{-4}$.
-Here we present results only for a single one-parameter family of charged initial data, which tells us nothing about universality.
+This corresponds to $|Q|/M\lesssim 10^{-4}$ for the black hole.
+So far we have presented results only for a single one-parameter family of charged initial data, which tells us nothing about universality.
 Both the mass and charge critical exponents appear to agree with their predicted values, however, which is indirect evidence of their universality.
 We also find that the fine-structure of the mass scaling law looks very similar in the charged and uncharged families.
-By contrast, we have found initial evidence that the fine-structure of the charge scaling law may not be universal.
-We will return to this elsewhere.
+By considering a more highly charged family, we have found initial evidence that the fine-structure of the charge scaling law may not be universal.
+We consider again complex Gaussian initial data with $x_{c}=1.5$, $d=0.5$, $\omega=2.0$ and ${\cal A}=0.05$, with initial gauge $G(0,x)=1$.
+The only difference is that we now set $q=10/\sqrt{4\pi}\simeq 2.821$, 10 times the value we investigated before.
+(As $q$ has dimension of inverse length, this is equivalent to keeping $q$ fixed and changing all other length scales in the numerical domain and initial data by factors of 10.) We find $p_{*}\simeq 1.157$.
+Once again we find charge scaling roughly with $\mu=0.883$, but the sign of the charge now oscillates in $\ln(p-p_{*})$.
+The charge fine-structure is shown in Fig.
+6.
+The mass scaling, by contrast, is unchanged (and therefore we do not show it here).
+We will return to this non-universality elsewhere.
+Figure 6: Charge fine-structure for a different family of complex Gaussian initial data (in the eG formulation, with $C_{\text{thr}}=1$.) The initial scalar field profile is the same, and again $G=1$ in the initial data, but $q$ is larger by a factor of 10, compared with the family represented in Fig.
+3.
+For plotting, we set A=0.23A=-0.23, otherwise as in Fig.
+3.
+The charge is positive where the curve is solid and negative where the curve is dashed.
+(The kinks in $\ln|Q|$ are where $Q$ goes through zero.) This is qualitatively different from Fig.
+3, where the charge is always positive.
 V.6 The role of the collapse criterion
@@
 From bottom to top, the curves correspond to $C_{\text{thr}}=0.8$ $0.9$, $0.95$, $0.99$, $0.999$ and $1$, as in Fig.
-7.
+8.
 The $C_{\text{thr}}=0.999$ and 11 curves are almost indistinguishable, but there is a tiny jump once per period in the $C_{\text{thr}}=0.999$ curve.
@@
 In the eG and fe formulations we can choose $C_{\text{thr}}=1$, but not in the eR formulation.
-In Paper II, for example, where we used the eR formulation, we used $C_{\text{thr}}=0.99$ in spherical symmetry and for mildly non-spherical initial data (and $C_{\text{thr}}=0.8$ for the most-nonspherical initial data), and found that the mass fine-structure was discontinuous, including in spherical symmetry.
+In Paper II, for example, where we used the eR formulation, we used $C_{\text{thr}}=0.99$ in spherical symmetry and for mildly non-spherical initial data, and $C_{\text{thr}}=0.8$ for the most-nonspherical initial data, and found that the mass fine-structure was discontinuous, including in spherical symmetry.
 To demonstrate the effect of the value of $C_{\text{thr}}$ on the measured $M$ and $Q$, Fig.
@@
 The exception is the location $x_{C}$ of the first occurrence of $C=1$: it is a very noisy function of $\ln(p/p_{*}-1)$, but with no obvious jumps.
-The noisiness is explained, in hindsight, by the difficulty of locating the first occurrence of $C=1$: in practice, we find the $x$-location where $C$ goes through 0 at the first timestep where its maximum is larger than 1.
+The noisiness is explained by the difficulty of locating the first occurrence of $C=1$: in practice, we find the $x$-location where $C$ goes through 0 at the first timestep where its maximum is larger than 1.
 Note that $M$ and $Q$ read off at this noisy location are not visibly noisy in Figs.
-2 and 3: we believe this is explained by the fact that $M$ and $Q$, as functions of $x$ along the curve $C=1$ quickly asymptote to $M_{\text{BH}}$ and $Q_{\text{BH}}$, and so the values of $M$ and $Q$ read off at $x=x_{C}$ depend only weakly on $x_{C}$.
-We can also understand theoretically why the mass fine-structure in spherical symmetry is continuous for $C_{\text{thr}}=1$ but discontinous for $C_{\text{thr}}<1$.
+2 and 3: we believe this is explained by the fact that $M$ and $Q$, as functions of $x$ along the curve $C=1$ (the apparent horizon) quickly asymptote to $M_{\text{BH}}$ and $Q_{\text{BH}}$, and so the values of $M$ and $Q$ read off at $x=x_{C}$ depend only weakly on $x_{C}$.
+We can understand theoretically why the mass fine-structure in spherical symmetry is continuous for $C_{\text{thr}}=1$ but discontinous for $C_{\text{thr}}<1$.
 We expect the function $C(u,x;p)$ to be continuous in all its arguments.
-Nevertheless, the set of points $C(u,x;p)=C_{\text{thr}}$ changes discontinously in $p$ whenever a local maximum of $C(u,x)$ rises above $C_{\text{thr}}$, and a new contour line forms.
-Therefore the location of the first 2-sphere where $C(u,x;p)=C_{\text{thr}}$ also depends discontinuously on $p$.
-$C_{\text{thr}}=1$ is an exception because the only contour line of $C(u,x;p)=1$ is the apparent horizon, dividing spacetime into a trapped and an untrapped region.
-But this exception does not apply beyond spherical symmetry, where a 2-sphere with $C=1$ is not in general marginally trapped.
-Based on our numerical experiments with different values of $C_{\text{thr}}$, and this theoretical understanding, we now believe that the discontinuity in the mass fine-structure observed in Paper II was an artifact of the choice $C_{\text{thr}}=0.99$, at least in spherical symmetry.
-Beyond spherical symmetry, we are not sure.
-We will return to this question elsewhere.
+Nevertheless, the set of points $(u,x)$ where $C(u,x;p)=C_{\text{thr}}$ is in general discontinuous in $p$, a simple example of catastrophe theory.
+The same is therefore true for the point in this set with the smallest value of $u$.
+However, $C_{\text{thr}}=1$ in spherical symmetry is an exception to this generic behaviour, because $C(u,x;p)=1$ is the apparent horizon, which divides spacetime into a trapped and an untrapped region, so it cannot jump in $p$.
+Beyond spherical symmetry, a 2-sphere with $C=1$ is not in general marginally trapped, and so does not lie in the apparent horizon.
+Therefore we generically expect the mass of the first coordinate 2-sphere with $C=1$ to jump in $p$.
+We will test this when we apply the eG formulation to axisymmetric critical collapse elsewhere.
 As a practical discovery, we find that plotting the equivalent of Fig.
-7 just for the data points obtained in bisection allows us to quickly find the appropriate value of $x_{0}$: if these curves drop to $x=0$ with increased fine-tuning, our choice of $x_{0}$ is too large (and we lose resolution in the self-similar region).
-If they rise to $x=x_{\text{max}}$ then $x_{0}$ is too small (and the first occurence of the specified threshold value of $C$ actually occurs at larger $x$, outside our numerical domain).
+8 just for the data points obtained in a bisection to $p_{*}$ allows us to quickly find the appropriate value of $x_{0}$: if these curves drop to $x=0$ with increased fine-tuning, our choice of $x_{0}$ is too large.
+If they rise to $x=x_{\text{max}}$ then $x_{0}$ is too small.
+To find $p_{*}$ and $x_{0}$, we then alternate an automated bisection in $p$ (to machine precision) with an adjustment of $x_{0}$ by hand (to about two to four digits).
 VI Tests on a null rectangle
@@
 That work uses double-null coordinates, compactified to both past and future null infinity (but finite across the past and future event horizon) as in MurReaTan13 .
-It uses the electromagnetic gauge $\nabla_{a}(R^{-2}A^{a})=0$, or $A_{u,v}+A_{v,u}=0$ in double-null coordinates.
-(This is in contrast to our electromagnetic gauge $A_{x}=0$, which in double-null coordinates becomes $A_{v}=0$.)
+It uses the electromagnetic gauge $\nabla_{a}(R^{-2}A^{a})=0$, or $A_{u,v}+A_{v,u}=0$ in double-null coordinates, in contrast to our electromagnetic gauge $A_{x}=0$, which in double-null coordinates becomes $A_{v}=0$.
 We are aware of only one previous numerical treatment of the full Einstein-charged scalar system in spherical symmetry: in BaakeRinne16 , the evolution is on hyperboloidal spacelike slices that on the left end are truncated with an excision boundary inside the event horizon of a RN black hole, and at their right end extend to future null infinity.
@@
 […]
-Our algorithm also allows us to fix the right spacetime gauge as $R_{,x}=R_{,x*}$ (so that $x$ is a linear funtion of $R$), but for brevity we do not present results here.
-When there is no scalar field on the right boundary, Bondi and affine right gauge are equivalent anyway.
+Our algorithm also allows us to fix the right spacetime gauge as $R_{,x}=R_{,x*}$, but for brevity we do not present results here.
+When there is no scalar field on the right boundary, Bondi and affine gauge are equivalent there anyway.
 Morever, in our discretisation they are equivalent without discretisation error.
-We can continue the spacetime gauge into the numerical domain with any choice of $B(u,x)$ (as long as $B=0$ on the left boundary), but again for brevity we only present double-null gauge $B(u,x)=0$, so that $x$ is an ingoing null coordinate everywhere.
-We use the timestep $\Delta u_{0}=C_{0}\Delta x$, with $C=0.1$.
+We can continue the spacetime gauge into the numerical domain with any choice of $B(u,x)$ (as long as $B=0$ on the left boundary), but again for brevity we only present double-null gauge $B(u,x)=0$, so that $x$ is an ingoing null coordinate everywhere, and the numerical domain is the full domain of dependence of the data (a null rectangle).
+We use the timestep $\Delta u_{0}=C_{0}\Delta x$, with $C_{0}=0.1$.
 We shall restrict to scenarios where $C_{*}<1$, so that the corner 2-sphere is untrapped.
 Then $R_{,x*}>0$ gives us $R_{,u*}<0$, and the ingoing Raychaudhuri equation (49) gives us $R_{,u}<0$ on the entire left boundary.
-We also restrict to the case where the right boundary is untrapped, so that $R_{,x}>0$ there.
+We also restrict to the case where the entire right boundary is untrapped, so that $R_{,x}>0$ there.
 VI.3 Prompt collapse
-Figure 8: Spacetime diagram of the prompt collapse scenario, showing the regular centre and numerical domain in bold, the singularity in bold-dashed, and the event horizon as a thin line.
-Figure 9: Spacetime diagram of collapse from data on a null cone with regular centre, with the lines as in Fig. 8. For suitable data, this could be the same spacetime as in Fig. 8.
+Figure 9: Spacetime diagram illustrating collapse from initial data on an outgoing null cone, with singularity excision.
+The regular centre and numerical domain are shown as bold lines, the singularity as bold dashed, lines of constant $u$ as thin solid, and lines of constant $x$ as thin dashed.
+The lines of constant $x$ illustrate sdn gauge, with $x=0$ the timelike regular centre $R=0$, $x=x_{0}$ ingoing null, and $x=x_{\text{max}}$ future spacelike.
+Initial data can be imposed on the outgoing null cone (with regular centre) $u=0$, symbolised here by a bell curve.
+Figure 10: Spacetime diagram illustrating the null rectangle setup, with initial data imposed on the ingoing null cone $x=0$ and/or the outgoing null cone $u=0$.
+The lines of constant $u$ (thin solid) are outgoing null again, as in Fig.
+9, but the lines of constant $x$ (thin dashed) are all ingoing null, illustrating double-null gauge.
+Figure 11: Spacetime diagram illustrating what we have called “prompt collapse”.
+Free initial data are imposed on $u=0$, but the data on $x=0$ are now specifically for Minkowski spacetime, so that the solution on the numerical domain can be continued into the triangle bounded by $R=0$, $x=0$ and $u=0$ (extended as thin solid lines) as a part of Minkowski spacetime.
 Consider a null rectangle with $R_{*}>0$, $R>0$ on the left boundary, $M_{*}=Q_{*}=0$, zero scalar field on the left boundary, and no trapped surface on the right boundary.
-Then the triangular region bounded by the regular centre $R=0$, $u=u_{0}$ and $v=v_{0}$ is a piece of Minkowski spacetime, and $u=u_{0}$ can be extended to the left to $R=0$ and to the right to $R=\infty$ as a lightcone with a regular vertex and no trapped surface.
+Then the triangular region bounded by the regular centre $R=0$, $u=u_{0}$ and $v=v_{0}$ is a piece of Minkowski spacetime, and $u=u_{0}$ can be extended to the left to $R=0$, with vanishing scalar field.
 If the data on $u=0$ are sufficiently strong that a MOTS first forms at $(v_{1},u_{\text{FMOTS}})$ for some $u_{0}<u_{\text{FMOTS}}<u_{1}$, then there must be an event horizon at some $u=u_{\text{EH}}$ with $u_{0}<u_{\text{EH}}\leq u_{\text{FMOTS}}$.
@@
 Figs.
-8 and 9 show gravitational collapse in the null rectangle and regular centre settings.
+9 and 10 show gravitational collapse from data on a single null cone, and from data on two intersecting null cones.
+Fig.
+11 illustrates the prompt collapse scenario.
+In each case, we illustrate excision of a future spacelike singularity from the numerical domain.
 If, on the other hand, there is no trapped surface in the null rectangle, formation of a black hole from the regular data on $u=u_{0}$ is neither proved nor disproved.
@@
 As an example of prompt collapse, we set ${\cal M}_{*}=Q_{*}=0$, $R_{*}=2$, and fix the corner gauge as $G_{*}=R_{,x*}=1$.
-With no scalar field on the left boundary, $R_{,u}=-1/2$ there, so the left boundary intersects $R=0$ at $u=4$.
+With no scalar field on the left boundary, $R_{,u}=R_{,u*}=-1/2$ there, so the left boundary intersects $R=0$ at $u=4$.
 On the right boundary we set $\psi$ to be a Gaussian with centre at $x_{c}=0.5$, width $d=0.1$, frequency $\omega=50$ and amplitude $9.5\cdot 10^{-3}$.
-We set $q=50$.
+We set $q=50$ and $x_{\text{max}}=1$.
 From now on, without loss of generality, we let $u_{0}=0$ and $v_{0}=v$, and we identify $v=0$ with $x=0$.
@@
 We set ${\cal M}_{*}=Q_{*}=1$, so that the horizon radius is again $r_{+}=1$, and we again set $R_{*}=2$.
-$M$ is now not constant, and in particular at the corner $M_{*}=3/4$, so that $C_{*}=3/8$.
+$M$ is now not constant, and in particular at the corner $M_{*}=3/4$, so that $C_{*}=3/4$.
 We again fix the gauge as $G_{*}=R_{,x*}=1$.
-This gives us $R_{,u}=-1/8$ on the left boundary, so that the event horizon is crossed at $u=8$, and the left boundary runs into the singularity at $u=16$.
+This gives us $R_{,u}=R_{,u*}=-1/8$ on the left boundary, so that the event horizon is crossed at $u=8$, and the left boundary runs into the singularity at $u=16$.
+We set $x_{\text{max}}=1$.
 On the horizon $R_{,x}=0$ but, in contrast to the subextremal case, inside the black hole $R_{,x}>0$ again, and the singularity $R=0$ is now timelike.
@@
 We evolve with $N_{x}=100\cdot 2^{n}$, with n=06n=0...6, up to $u=7.9$ in the eR formulation, and up to $u=14$ in the eG and fe formulations.
-We see perfect self-convergence and convergence between formulations in the highest resolutions, but convergence in the lower resolutions begins to break down between $u=13$ and $u=14$ due to the appearance of sharp gradients in all variables near the left boundary.
+We see perfect self-convergence and convergence between formulations in the highest resolutions, but convergence in the lower resolutions begins to break down between $u=13$ and $u=14$, due to the appearance of sharp gradients in all variables near the left boundary.
 VI.5 Perturbed extremal Reissner-Nordström
@@
 Not unexpectedly, this particular attempt fails.
-Indeed for all $u>1.7$, already well before the putative horizon is crossed, $Q\leq{\cal M}\leq 1$ everywhere on the domain $0\leq x\leq 1$ (with equality only at the left boundary, where the solution is still electrovacuum extremal RN).
-The event horizon is crossed at $u\simeq 9$ and has radius $R\simeq 0.9$ (and hence Hawking mass $M\simeq 0.45$), augmented mass ${\cal M}\simeq 0.996$ and charge $Q\simeq 0.991$, so the final black hole is slightly subcritical.
-However, while $R$ on our numerical domain $0\leq x\leq 1$ levels off in $x$ at $u\simeq 9$, it returns to being an increasing function of $x$ afterwards, at least up to $u=14$, when we stop the simulation.
-Although this behaviour is characteristic of an extremal black hole (which has no trapped surfaces in the interior), we know in this case that the black hole is clearly subcritical.
-Hence the fact that we do not see $R_{,x}<0$ or the spacelike singularity $R=0$ on our numerical domain must be due to its limited size.
-We evolve with $N_{x}=100\cdot 2^{n}$, now with n=06n=0...6.
+Indeed for all $u>1.7$, already well before the putative horizon is crossed, $Q\leq{\cal M}\leq 1$ everywhere on the domain $0\leq x\leq 1$ (with equality only at the left boundary).
+The event horizon is crossed at $u\simeq 7.4$, that is, earlier than in the unperturbed eRN solution and has radius $R\simeq 1.08$ (and hence Hawking mass $M\simeq 0.54$), augmented mass ${\cal M}\simeq 0.997$ and charge $Q\simeq 0.993$, so the final black hole is slightly subcritical.
+$R$ on our numerical domain $0\leq x\leq 1$ levels off in $x$ at $u\simeq 7.4$, indicating the event horizon, then has a maximum in $x$ between $u\simeq 7.4$ and $u\simeq 8.8$, and then returns to be an increasing function of $x$ at least up to $u=14$, when we stop the simulation.
+This is illustrated in Fig.
+12.
+That we do not see $R_{,x}<0$ for sufficiently large $x$, for all $u>u_{\text{horizon}}$, is probably due to the limited size of our numerical domain.
+We will investigate this elsewhere.
+Figure 12: Plot of the area radius $R(u,x,)$ in the eRN solution perturbed by a complex Gaussian on the right boundary $u=0$.
+We plot against $x$, at $u=0$, $7.2$, $7.4$, $7.6$ and $14$ (from top to bottom).
+The lower plot shows only $u=7.2$, $7.4$ and $7.6$ for clarity, demonstrating that the event horizon is crossed at $u\simeq 7.4$.
+We evolve with $N_{x}=100\cdot 2^{n}$, with n=06n=0...6.
 We need the higher resolutions to see clean self-convergence and convergence between resolutions.
-The simulation is more demanding than that of unperturbed eRN because $\phi$ becomes more rapidly oscillating in $x$: in the initial data, $\phi$ has a about two complex oscillations within a Gaussian envelope, while at $u=14$ there are about 30 complex oscillations in roughly the same envelope.
+The simulation is more demanding than that of unperturbed eRN because $\phi$ becomes rapidly oscillating in $x$.
+This is illustrated in Fig.
+13.
+Figure 13: Plot of the real part $\psi(u,x,)$ of the complex scalar field in the eRN solution perturbed by a complex Gaussian on the right boundary $u=0$.
+We plot against $x$, at $u=0$ (the simpler graph) and at $u=14$ (the highly oscillating graph).
+The imaginary part $\chi$ looks similar, with a phase shifting indicating that the rapid oscillation is in the complex phase.
+The numerical domain is $0\leq x\leq 1$, but we plot only the range where $\psi$ is large.
 […]
 Once again we see good convergence within formulations and between formulations at the highest resolution.
-At $u=14$, the complex scalar field is rapidly oscillating in $x$, with about 20 phase oscillations between $x=0$ and $x=0.1$, with an exponentially decaying envelope, and this is only just resolved at the two highest resolutions.
+At $u=14$, the complex scalar field is rapidly oscillating in $x$, with about 16 phase oscillations between $x=0$ and $x=0.1$, with an exponentially decaying envelope, and this is only just resolved at the two highest resolutions.
 VII Conclusions
-This series of papers investigates the usefulness of “Bondi-like” null coordinates, in which surfaces of constant retarded time $u$ are null cones and the coordinate lines of constant $(u,\theta,\varphi)$ are their generators.
+This series of papers investigates the usefulness of single null coordinates, in which surfaces of constant retarded time $u$ are null cones and the coordinate lines of constant $(u,\theta,\varphi)$ are their generators.
 The remaining gauge freedom is how points on each generator are labelled by the radial coordinate $x$.
@@
 There are also two particular formulations each tied to a specific gauge: the Bondi formulation where $R=x$, and the affine formulation where $G=1$.
+These are maximally constrained, in the sense that only physical degrees of freedom are evolved.
 To our knowledge, there is no clear statement in the literature that any of the three general formulations can be combined with any choice of radial gauge.
 This observation is the main conceptual novelty of the present paper.
-To present the five different formulations pedagogically, and with a focus on regularity at the central worldline $R=0$, in this paper we have restricted attention to spherical symmetry.
-We will explore the eG formulation in axisymmetry in a future paper, but we notice already that the generalisation beyond spherical symmetry and/or to higher dimensions is conceptually straightforward.
-As a nontrivial test-bed in spherical symmetry, we have implemented the spherical Einstein-Maxwell-charged scalar system, allowing for initial data either on a single outgoing null cone $u=0$ emerging from a regular centre $R=0$, or on two intersecting null cones $u=u_{0}$ and $v=v_{0}$.
+The main motivation for introducing the eG and fe formulations, in which $G$ is evolved, is that in them we can evolve through event and apparent horizons.
+In Paper II, both our choice of the eR formulation and our choice of lsB gauge did not allow us to do this in axisymmetric scalar collapse.
+This in turn meant that we could not bisect arbitrarily close to the collapse threshold in highly-nonspherical one-parameter families of initial data.
+In the eG and fe formulations, we can now evolve through horizons, while still using $B$ to make the coordinate adjust to the approximate self-similarity observed in type-II critical collapse.
+We plan to apply our eG formulation with a suitable radiable gauge to this problem.
+To present the new formulations pedagogically, and with a focus on regularity at the central worldline $R=0$, in this paper we have restricted attention to spherical symmetry.
+The generalisation beyond spherical symmetry and/or to higher dimensions is conceptually straightforward, and will be given elsewhere.
+As an example in spherical symmetry that is also of interest in its own right, we have implemented the spherical Einstein-Maxwell-charged scalar system.
+We allow for initial data either on a single outgoing null cone $u=0$ emerging from a regular centre $R=0$, or on two intersecting null cones $u=u_{0}$ and $v=v_{0}$.
 In both setups we have shown second-order self-convergence in all three formulations and second-order convergence between all three formulations.
-With a regular centre, the convergence is pointwise except near the centre, where it still holds in any norm, such as the root-mean-square or maximum norm.
+With a regular centre, the convergence is pointwise except near the centre, where it still holds in any function norm.
 The fact that even the fe formulation converges without any tweaks at the centre came as a pleasant surprise.
 With the left boundary an ingoing null cone, the convergence is perfectly pointwise everywhere.
-As a physics testbed of collapse from regular data on an outgoing null cone, we have investigated critical collapse of a charged scalar field.
-This had previously been investigated theoretically in GundlachMartin96 , and numerically in HodPiran97 ; Petryk05 , but here we show the first accurate plots of the universal fine structure in the mass and charge scaling laws.
-Surprisingly, we found that being able to evaluate the mass and charge on a marginally trapped surface makes a qualititative difference to our results, compared to evaluating them on an almost-trapped surface, even one with Hawking compactness as large as $C=0.999$.
-We also found initial evidence that the fine-structure of the charge scaling law may not be universal.
-We will return to this elsewhere.
-However, the main motivation for introducing the two formulations in which $G$ is evolved is that in them we can evolve through event and apparent horizons.
-In Paper II, both our formulation (we evolved $R$) and our choice of gauge (“local shifted Bondi gauge”) did not allow us to do this in axisymmetric scalar collapse.
-This in turn meant that we could not bisect arbitrarily close to the collapse threshold in highly-nonspherical one-parameter families of initial data.
-We plan to apply our eG formulation with a suitable radiable gauge to this problem.
-As a physics testbed of the setup with initial data on two intersecting null cones, we have simulated gravitational collapse with a Minkowski interior, the exact Schwarzschild and extreme Reissner-Nordström solutions (which are non-trivial in our gauge), and extreme Reissner-Nordström perturbed by a complex scalar field on either $u=u_{0}$ or $v=v_{0}$.
-We plan a study of the solution space of the spherical Einstein-Maxwell-charged scalar system, and in particular an investigation of the “extremal critical collapse conjecture for the spherical Einstein-charged scalar field system KehleUnger24 , see also KehleUnger22 .
+As a physics testbed of collapse with data on an outgoing null cone, we have investigated spherical type-II critical collapse of a charged scalar field.
+This had previously been investigated theoretically in GundlachMartin96 , and numerically in HodPiran97 ; Petryk05 , but here we have presented the first accurate plots of the universal fine structure in the mass and charge scaling laws.
+We have also presented some initial evidence for non-universality of the charge scaling fine-structure.
+As testbeds for the setup with initial data on two intersecting null cones, we have simulated gravitational collapse with a Minkowski interior, the exact Schwarzschild and extreme Reissner-Nordström solutions (which are non-trivial in our gauge), and extreme Reissner-Nordström perturbed by a complex scalar field on either $u=u_{0}$ or $v=v_{0}$.
+We plan a study of the solution space of the spherical Einstein-Maxwell-charged scalar system, and in particular an investigation of the “extremal critical collapse conjecture” for the spherical Einstein-charged scalar system KehleUnger24 , see also KehleUnger22 .
 We anticipate that both types of numerical domain will be required for this investigation.
@@
 $f(r):=1-{2{\cal M}_{0}\over r}+{Q_{0}^{2}\over r^{2}},$ (86)
-and the real constants ${\cal M}_{0}>0$ and $Q_{0}$ are the usual mass and charge parameters of the RN metric, normalised such that $|Q_{0}|={\cal M}_{0}$ represents an extremal black hole.
-The event and Cauchy horizon are at
+[…]
 $r_{\pm}={\cal M}\pm\sqrt{{\cal M}^{2}-Q^{2}}.$ (87)
@@
 $M=r_{+}/2$ at the event horizon is also the irreducible mass of the black hole.
-In the special case of an extremal black hole, $r_{+}=r_{-}=2M={\cal M}=Q$ on the event horizon.
+In the special case of an extremal black hole, $r_{+}=r_{-}=2M={\cal M}=Q$.
 The tortoise radius $r_{*}$ is defined, up to addition of a constant, by
@@
 $B$ $=$ ${1\over 2}\left(1-{2{\cal M}_{0}\over x}+{Q_{0}^{2}\over x^{2}}\right).$ (95)
-[…]
+In the double-null coordinates $u=\bar{u}$, $x=\bar{v}$ we have
 $R$ $=$ $r\left({\bar{v}-\bar{u}}\over 2\right),$ (96)
@@
 $r_{0}:=r\left({\bar{v}_{0}-\bar{u}_{0}\over 2}\right),$ (101)
-we have $u(\bar{u}_{0})=u_{0}$, $v(\bar{v}_{0})=v_{0}$, $R_{,u}=-1/a$ along $v=v_{0}$ and $R_{,v}=1/b$ along $u=u_{0}$. Inverting, we obtain
+we have $u(\bar{u}_{0})=u_{0}$, $v(\bar{v}_{0})=v_{0}$, $R_{,u}(u,v_{0})=R_{,u*}:=-1/a$ and $R_{,v}(u_{0},v)=(R_{,v})_{*}:=1/b$. Inverting (99,100), we obtain
 $\bar{u}(u)$ $=$ $\bar{v}_{0}-2r_{*}\left(r_{0}+{u-u_{0}\over-a}\right),$ (102)
@@
 $G(u,v)={\bar{G}(\bar{v},\bar{u})\over ab\,\bar{G}(\bar{v}_{0},\bar{u})\,\bar{G}(\bar{v},\bar{u}_{0})},$ (107)
-and hence $G(u_{0},v)=G(u,v_{0})=G_{*}$. This is, in implicit form, the exterior metric of RN in the coordinates we use in our null-rectangle tests, with $(R_{,v})_{*}=1/b$, $R_{,u*}=-1/a$, and
-$G_{*}={1\over ab\,\bar{G}(\bar{u}_{0},\bar{v}_{0})}.$ (108)
+and hence $G(u_{0},v)=G(u,v_{0})=G_{*}$, where
+$G_{*}:={1\over ab\,\bar{G}(\bar{u}_{0},\bar{v}_{0})}.$ (108)
+This is, in implicit form, the exterior metric of RN in the coordinates we use in our null-rectangle tests.
 Appendix B Regular centre
-The Minkowski metric in Bondi coordinates, with $U$ normalised to be proper time at the centre, is
+The Minkowski metric in Bondi coordinates, with $U$ the retarded time normalised to be proper time at the centre, is
 $ds^{2}=-dU^{2}-2dU\,dR+R^{2}\,d\Omega^{2}.$ (109)
@@
 ${\cal H}(u,0)$ $=$ $0.$ (117)
-From elementary flatness, all these conditions at the centre must also hold in an arbitary curved spherical spacetime with a regular centre $R=x=0$ and $u$ proper time at the centre.
-Beyond spherical symmetry, in spherical coordinates, they will also hold at the regular “central” worldline $R=x=0$ as long as it is geodesic, see Appendix III.E.2 of Paper I for this subtlety.
+From elementary flatness, these conditions at the centre must also hold in an arbitary curved spherical spacetime with a regular centre $R=x=0$ and $u$ proper time at the centre.
+Beyond spherical symmetry, in spherical coordinates, they will also hold at the regular worldline $R=x=0$ as long as it is geodesic, see Appendix III.E.2 of Paper I.
 Appendix C Expansions at the centre
 For finding $G$ by integration, given $R$, we need
-$\ln\left({G\over R_{,x}}\right)$ $=$ $2\pi(\psi_{(1)}^{2}+\chi_{(1)}^{2})x^{2}$ (118)
+$\ln\left({G\over R_{,x}}\right)$ $=$ $2\pi(\psi_{(1)}^{2}+\chi_{(1)}^{2})x^{2}$
 $+{4\pi\over 3}\Bigl[4(\psi_{(1)}\psi_{(2)}+\chi_{(1)}\chi_{(2)})$
-$-{R_{(2)}\over R_{(1)}}(\psi_{(1)}^{2}+\psi_{(2)}^{2})\Bigr]x^{3}+O(x^{4}).$
+$-{R_{(2)}\over R_{(1)}}(\psi_{(1)}^{2}+\psi_{(2)}^{2})\Bigr]x^{3}$
+$+O(x^{4}).$ (118)
 Equivalently, for finding $R$ by solving a system of two first-order ODEs, given $G$, we require
@@
 $+{R_{(1)}\over 2}\Bigl[R_{(2)}(\psi_{(0)}\chi_{(1)}-\chi_{(0)}\psi_{(1)})$
-$+R_{(1)}(\psi_{(0)}\chi_{(2)}-\chi_{(0)}\psi_{(2)})\Bigr]x^{4}+O(x^{5}),$
+$+R_{(1)}(\psi_{(0)}\chi_{(2)}-\chi_{(0)}\psi_{(2)})\Bigr]x^{4}$
+$+O(x^{5}),$
 ${A\over{4\pi}q}$ $=$ ${R_{(1)}\over 6}(\chi_{(1)}\psi_{(0)}-\psi_{(1)}\chi_{(0)})x^{2}$ (123)
 $+{1\over 6}\Bigl[R_{(2)}(\psi_{(0)}\chi_{(1)}-\chi_{(0)}\psi_{(1)})$
-$+R_{(1)}(\psi_{(0)}\chi_{(2)}-\chi_{(0)}\psi_{(2)})\Bigr]x^{3}+O(x^{4}).$
+$+R_{(1)}(\psi_{(0)}\chi_{(2)}-\chi_{(0)}\psi_{(2)})\Bigr]x^{3}$
+$+O(x^{4}).$
 Note that $R_{(3)}$ (or $G_{(2)}$) does not appear to this order.
@@
 $-{4\pi\over 3}\Biggl[{R_{(2)}\over R_{(1)}^{2}}(\psi_{(1)}^{2}+\chi_{(1)}^{2})$
-$-{3\over R_{(1)}}(\psi_{(1)}\psi_{(2)}+\chi_{(1)}\chi_{(2)})\Biggr]x^{2}+O(x^{3}).$
+$-{3\over R_{(1)}}(\psi_{(1)}\psi_{(2)}+\chi_{(1)}\chi_{(2)})\Biggr]x^{2}$
+$+O(x^{3}).$
 Appendix D Radial null geodesics peeling off the horizon
@@
 We note that on RN $\bar{v}$ obeys $\bar{v}=\bar{u}+2R$.
-If we assume that for sufficiently early $u$ and sufficiently late $\bar{v}$ the collapse spacetime is approximated by the RN exterior, we can characterize $\bar{v}$ on the collapse spacetime as the retarded time that obeys $\left.R_{,\bar{v}}\right|_{u}=2$ in that region.
-We could then define $\bar{v}=2R$ on $u=0$ and propagate it with $\Xi\bar{v}=0$.
+If we assume that for sufficiently early $u$ and sufficiently late $\bar{v}$ the collapse spacetime is approximated by the RN exterior, we can define $\bar{v}=2R$ on $u=0$ and propagate it with $\Xi\bar{v}=0$.
 Appendix E Boundary conditions on an ingoing null cone intersecting an outgoing one
@@
 The ingoing Raychaudhuri equation (48) with $A=B={\cal H}=0$ gives
-$\ddot{R}_{l}+4\pi R_{l}(\dot{\psi}_{l}^{2}+\dot{\chi}_{l}^{2})=0.$ (133)
+$\ddot{R}_{l}+4\pi(\dot{\psi}_{l}^{2}+\dot{\chi}_{l}^{2})R_{l}=0.$ (133)
 Note that $\ddot{R}_{l}\leq 0$, so if $\dot{R}_{l*}<0$ at the corner (as we shall always assume), then $\dot{R}_{l}<0$ for all $u\geq 0$. Eq. (28) with $A=B=0$ gives
@@
 ${2\dot{R}_{l}R_{,xl}\over G_{l}}=C_{l}-1,$ (135)
-where $C_{l}$ is used as a shorthand for
+where $R_{,xl}(u):=R_{,x}(u,0)$ and $C_{l}$ is used as a shorthand for
 $C_{l}:={2\over R_{l}}\left({\cal M}_{l}-{Q_{l}^{2}\over 2R_{l}}\right),$ (136)
@@
 $V_{l}={C_{l}-1\over 2\dot{R}_{l}}.$ (139)
-We now have the data that we need to start up the integration of the hierarchy equations at $x=0$, namely $Q=Q_{l}$ for (27), $A=0$ for (30), either $R=R_{l}$ and $V=V_{l}$ for (64,65) or $G=G_{*}$ for (39), $\Xi R=\dot{R}_{l}$ for (42), $\hat{\Xi}\psi=\dot{\psi}_{l}$ for (43), $\hat{\Xi}\chi=\dot{\chi}_{l}$ for (44), and ${\cal H}=0$ for (45).
+We now have the data that we need to start up the integration of the hierarchy equations at $x=0$, namely $Q=Q_{l}$ for (27), $A=0$ for (30), either $R=R_{l}$ and $V=V_{l}$ for (65,64) or $G=G_{*}$ for (39), $\Xi R=\dot{R}_{l}$ for (42), $\hat{\Xi}\psi=\dot{\psi}_{l}$ for (43), $\hat{\Xi}\chi=\dot{\chi}_{l}$ for (44), and ${\cal H}=0$ for (45).
 In the simple case where $\psi_{l}=\chi_{l}=0$ the boundary ODE system becomes trivial, giving $Q_{l}=Q_{*}$, ${\cal M}_{l}={\cal M}_{*}$, and $R_{l}=R_{*}+\dot{R}_{l*}\,u$.
@@
 With $A=B=0$ on the left boundary, the values of $\psi_{,u}$, $\chi_{,u}$, $R_{,u}$ and/or $(\ln G)_{,u}=0$ are then the same, up to round-off error, in the boundary ODE system and at $x=0$ in the bulk PDE system.
-The Runge-Kutta ODE solver that we use for the time integration assumes that the ODE system is autonomous, we add $u$ as an evolved variable to the set of left-boundary variables, in order to access the given functions $\dot{\psi}_{l}(u)$ and $\dot{\chi}_{l}(u)$ at the correct substep values of $u$.
+The Runge-Kutta ODE solver that we use for the time integration assumes that the ODE system is autonomous.
+We add $u$ as an evolved variable to the set of left-boundary variables, in order to access the given functions $\dot{\psi}_{l}(u)$ and $\dot{\chi}_{l}(u)$ at the correct substep values of $u$.
 Appendix G Convergence testing
@@
 For a numerical quantity at resolution $h$, given that our numerical methods are designed to be second-order accurate, we expect the Richardson expansion
-$f(u,x;h)=f_{0}(u,x)+h^{2}f_{2}(u,x)+h^{3}f_{3}(u,x)+O(h^{4}),$ (141)
-where $f_{0}(u,x)$ is the (unknown) continuum solution and $f_{i}(u,x)$ are (discretisation-dependent, unknown) error terms, and hence
+$f(u,x;h)=f_{0}(u,x)+h^{2}f_{2}(u,x)+h^{3}f_{3}(u,x)+O(h^{4})$ (141)
+to hold, where $f_{0}(u,x)$ is the (unknown) continuum solution and $f_{i}(u,x)$ are (discretisation-dependent, unknown) error terms, and hence
 ${\cal E}[f](u,x;h)=h_{\text{ref}}^{2}\left({3\over 4}f_{2}(u,x)+{7\over 8}hf_{3}(u,x)+O(h^{2})\right).$ (142)